### v1

# Driving Scene Synthesis on Free-form Trajectories with Generative Prior

Abstract
Driving scene synthesis along free-form trajectories is essential for driving simulations to enable closed-loop evaluation of end-to-end driving policies. While existing methods excel at novel view synthesis on recorded trajectories, they face challenges with novel trajectories due to limited views of driving videos and the vastness of driving environments. To tackle this challenge, we propose a novel free-form driving view synthesis approach, dubbed DriveX, by leveraging video generative prior to optimize a 3D model across a variety of trajectories. Concretely, we crafted an inverse problem that enables a video diffusion model to be utilized as a prior for many-trajectory optimization of a parametric 3D model (e.g., Gaussian splatting). To seamlessly use the generative prior, we iteratively conduct this process during optimization. Our resulting model can produce high-fidelity virtual driving environments outside the recorded trajectory, enabling free-form trajectory driving simulation. Beyond real driving scenes, DriveX can also be utilized to simulate virtual driving worlds from AI-generated videos.
Figure 1: Though conventional reconstruction method can well fit views on the recorded trajectory, they struggle to extrapolate onto the views along the novel trajectories (top). In contrast, our method significantly improves the synthesis quality on these novel views (bottom), thereby achieving driving scene synthesis on free-form trajectories.
1 Introduction
Building virtual driving worlds engagible with driving policies plays a pivotal role in developing robust autonomous driving systems. It enables automatic synthesis of diverse driving data, including safety-critical long-tail scenarios, edge cases, and providing closed-loop evaluation for end-to-end autonomous driving systems.
Recent advancements in novel view synthesis [21, 15] have driven reconstruction-based driving simulation [34, 4, 35] due to its ability to synthesize photorealistic sensor data. These methods aim to reconstruct driving environments from individual driving sequences, typically involving single-trajectory videos from surrounding cameras with sparse views, minimal overlap, and textureless areas. Despite impressive results in fitting training views and interpolating along original trajectories, they still struggle to extrapolate substantially novel views far from the recorded trajectory, leading to limited flexibility and usefulness for driving simulation. To address these problems, a couple of latest works [13, 38] explored image generative prior as extra novel view regularization. However, they are largely limited in design due to (i) the need for fine-tuning an off-the-shelf image generation models to obtain the capability of view transformation, suffering from both additional computational cost and per-scene data scarcity, and (ii) absence of the desired spatiotemporal knowledge within image models. To tackle these issues, a concurrent work [40] uses a pretrained video diffusion model to generate novel trajectory training images conditioned on initial recorded frames. Without sufficient content constraint but with intrinsic hallucination risk, this approach however tends to drift away from the original scene in scene generation, making troubles for the subsequent model optimization.
To address the challenges mentioned above, we propose a novel driving scene synthesis approach for free-form trajectories, termed DriveX, by leveraging video generative prior with rich spatiotemporal knowledge during the model (e.g., Gaussian splatting) optimization. This is achieved by designing an inverse problem that enables a video diffusion model to serve as a prior for optimizing the Gaussian over a variety of trajectories. During training, we render novel trajectory views using the in-training model. To identify potential artifacts from trajectory extrapolation, we compare each rendered view with recorded images wrapped under the same view and geometric conditions, allowing us to obtain reliable regions for each rendered image. The inverse problem is then crafted as restoring each rendered view via utilizing its corresponding reliable regions as conditioning inputs, which can be effectively executed by the video diffusion model. The output from this process is subsequently fed into the optimization as supervision to update the target model. To seamlessly exploit the generative prior, we iteratively conduct this process during Gaussian optimization. As shown in Fig. 1, our approach substantially improve the view synthesis quality in novel trajectory compared to the traditional reconstruction methods.
The contributions of this paper are as follows: (i) We propose leveraging video generative prior with rich spatiotemporal knowledge for generalizable driving scene synthesis from single-trajectory recorded videos. (ii) We introduce DriveX, a novel driving scene synthesis framework for free-form trajectories that innovatively constructs an inverse problem to enable the use of a video diffusion model as a prior. (iii) Extensive experiments demonstrate that DriveX significantly outperforms existing state-of-the-art alternatives in driving scene synthesis using single-trajectory recorded videos. Additionally, we showcase the superiority of our method in rendering novel trajectories of virtual driving worlds from AI-generated videos, even when faced with inherent content inconsistencies. This approach enables a more economically scalable large-scale simulation without the need to collect videos for each scene.
2 Related works
Reconstruction-based driving scene synthesis
Early approaches [34, 36, 27, 32, 28] for dynamic urban scenes reconstruction focus on extending neural radiance fields (NeRF) [21] in large scale unbounded scenes.These methods are restricted with low efficiency and blurry rendering performance. Recently, another line of works introduced 3D Gaussian splatting [15] into this task to produce real-time high-fidelity rendering. Some of them [35, 42] leverage its favorable explicit property to decouple the whole scene into the static background and moving foreground Gaussians, while others modeling the motion of 3D Gaussians with periodic functions [4] or deformation fields [12]. However, driving scenes featured with sparse views, little overlaps, and large textureless areas, making existing methods struggle to generalize to substantively novel views.
Generative-based driving scene synthesis
Recently, propelled by the significant advancement of the diffusion model [10, 24, 22, 25, 1], many works employ video generators to construct world models [16]. Some of them introduce this paradigm into autonomous driving [29, 41, 17, 6, 30, 7, 11, 20] for simulating the plausible visual outcomes under versatile driving controls. Despite the remarkable diversity of generated results, this approach is constrained by the absence of underlying 3D model. Consequently, they cannot promise geometry and texture consistency for the generated many-trajectory videos of the same scene.
Figure 2: The schematic illustration of DriveX. Top: Gaussian optimization with both recorded trajectory and generated novel trajectory as supervision. Bottom: Generative prior integration. Given a novel view rendered image from Gaussians $\mathcal{G}_{t}$, video diffusion is applied to generate its refined version with condition of unreliability mask. We omit the loss on recorded trajectory for clarity.
Scene reconstruction with diffusion prior
Leveraging generative prior to enhance the reconstruction from sparse view observations [37, 19, 18] has emerged. Extending this paradigm into driving scene reconstruction has been recently attempted [38, 13, 40]. However, they typically suffer from limited diversity and scale of driving scene data. SGD [38] and VEGS [13] contribute to define and tackle pseudo and extrapolated view synthesis in driving scenarios. However, these methods may encounter challenges with accurate pose and detail estimation under novel trajectories. GGS [8] introduces virtual lane generation module, but its dependency on modified MVSplat structure poses challenges for generalization to large-scale pretraining. A concurrent work [40] uses a world model [29] to generate novel trajectory video referenced on initial frame for trajectory augmentation. However, it tends to drift away from the original scene due to the strong built-in hallucination risk and insufficient constraints. This would inevitably hurt the subsequent model optimization. In contrast, our method dedicates to solve an inverse problem that enables a video diffusion model to be used as a prior for many-trajectory optimization of a parametric 3D model, achieving superior synthesis quality on free-form trajectories.
3 Methodology
3.1 Preliminary: Driving scene reconstruction
Reconstruction-based driving simulation aims to recover the surrounding scenes from calibrated video $\mathcal{V}_{\mathrm{gt}}=\left\{I_{i}\right\}_{i=0}^{F}$ recorded along a driving trajectory $\mathcal{T}=\left\{P_{i}\right\}_{i=0}^{F}$, where $P_{i}\in\mathrm{SE(3)}$ is the camera pose at frame $i$. Then the sensor data would be rendered when the vehicle performs simulated unrecorded actions (e.g. lane change) in the reconstructed environment.
Recently, due to the superior fidelity and efficiency of 3D Gaussian splatting (3DGS) [15], it has been widely employed in numerous street scene reconstruction efforts [35, 42, 4, 12] as a fundamental scene representation. 3DGS represents scenes as a set of 3D Gaussians, denoted $\mathcal{G}$. The influences of each Gaussian primitive with position $\mathbf{x}_{k}\in\mathbb{R}^{3}$ and covariance $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ on any position is given by an unnormalized Gaussian function $G_{k}(\cdot;\mathbf{x}_{k},\mathbf{\Sigma})$ weighted by its opacity $o_{k}\in\mathbb{R}$. Given any camera pose $P_{i}$, it can be rendered by performing splatting and alpha blending on $N$ sorted Gaussians visible at this view:
$I^{\prime}=R(\mathcal{G},P_{i})=\sum^{N}_{k=1}o_{k}G_{k}c_{k}(P_{i})\prod^{k-1}_{j=1}(1-o_{j}G_{j}),$ (1)
where $R$ represents the differentiable Gaussian rasterizer, $I^{\prime}$ denotes the rendered image, $c_{k}$ is the view-dependent color of the $k$-th Gaussian. Then the Gaussians $\mathcal{G}$ can be optimized via the photometric loss between rendered images $I_{i}^{\prime}$ and ground truth $I_{i}$ from recorded video $\mathcal{V}_{\mathrm{gt}}$ by:
$\mathcal{L}_{\text{img}}(I_{i}^{\prime},I_{i})=\lambda\|I_{i}^{\prime}-I_{i}\|_{1}+(1-\lambda)\mathcal{L}_{\text{SSIM}}(I_{i}^{\prime},I_{i}),$ (2)
where $\mathcal{L}_{\text{SSIM}}$ is the SSIM [31] loss and $\lambda$ is a weight.
In the urban scenes, although the traffic participants are dynamic, they can be regarded as static in their own local coordinate frames. The transformation from each object’s local frame to the global coordinate frame of the whole scene is parameterized by a series of rotations and transformations, which can be derived from its tracklet annotated by off-the-shelf 3D trackers [5, 33]. Then a complete dynamic street scene can be modeled by decoupled foreground moving Gaussians and static background Gaussians. Additionally, cube map is commonly used to model the sky with infinite distance [35, 4].
Challenges
Since the camera configuration in autonomous vehicles is primarily designed for perception tasks, the captured videos often offer limited viewpoints, minimal overlap, and large textureless regions. These inherent properties bring challenges to the reconstruction of the unbounded driving scenes from single-trajectory videos. Fitting the training views at recorded trajectory $\mathcal{T}$ is insufficient to realize satisfactory extrapolation out of $\mathcal{T}$ for novel view synthesis. As a result, existing optimization-based driving scene synthesis approaches struggle to achieve high-quality view rendering along novel trajectories $\mathcal{T}^{\prime}=\{P^{\prime}_{i}|P^{\prime}_{i}\notin\mathcal{T}\}_{i=0}^{F}$, often leading to obvious degeneration and artifacts.
3.2 DriveX
To address the aforementioned challenges, we propose a novel framework, DriveX, seamlessly integrating generative priors into the driving scene reconstruction process (e.g. Gaussian optimization). As shown in Fig. 2, it iteratively synthesizes novel trajectory scenes by utilizing a video diffusion as supervision to optimize the Gaussian model. To ensure the generated content aligns with the underlying scenes in detail, the generation process is appropriately conditioned on the in-training Gaussian model.
To achieve this, we delicately design an inverse problem [14, 23], which can be formulated as recovering clean novel view images $\mathcal{V}$ from the artifact-heavy novel view rendered images $\mathcal{V}^{\prime}$ on novel trajectories $\mathcal{T}^{\prime}$. Concretely, $\mathcal{V}^{\prime}$ can be regarded as a measurement of unknown ground truth $\mathcal{V}$, as:
$\mathcal{V}^{\prime}=f(\mathcal{V})+\epsilon,$ (3)
where $\epsilon$ is a random noise, and the measurement function $f$ distorts $\mathcal{V}$ due to poor generalization ability of Gaussian model in conventional methods. With this design, a video generative model [37] can be readily employed to address this problem effectively. The refined video can then serve as additional supervision within Gaussian optimization.
Solving inverse problem with generative prior
The key to solve the inverse problem lies in accurately identifying the artifacts in the novel view rendered video. Inspired from the sparse view reconstruction methods [34, 19], we construct an unreliability mask representing the artifacts via comparing the rendered images with recorded images, which is used to directly guide the diffusion model focus on refining these unreliable regions while preserving reliable areas.
Specifically, given a sequence of viewpoints $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ along a novel trajectory, a video diffusion model $\mathcal{D}$ is employed to estimate $\mathcal{V}_{i}$ in Eq. 3:
$\mathcal{V}^{\prime}_{t}$ =$=\left\{R(\mathcal{G}_{t},~{}P_{i}^{\prime})\right\}_{i=0}^{F},$ (4)
$\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ $=\mathcal{D}\left(\mathcal{V}^{\prime}_{t},~{}\mathcal{M}\right),$ (5)
$\mathcal{L}^{\prime}$ $\triangleq\mathcal{L}_{\text{img}}(\mathcal{V}^{\prime}_{t},~{}\mathcal{V}^{\prime}_{t,\mathrm{refine}}),$ (6)
where $\mathcal{V}^{\prime}_{t}$ is rendered from the Gaussian model $\mathcal{G}_{t}$ at current iteration $t$ through Eq. 1, the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ can be regard the estimation of $\mathcal{V}$ in Eq. 3, and $\mathcal{M}$ is the mask indicating unreliable regions. Typically, $\mathcal{D}$ takes noise-perturbed images as input, conditioned on the masked video, and outputs the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$. Then $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ is used as the additional supervision for the novel trajectory loss $\mathcal{L}^{\prime}$ (Eq. 6) to improve the novel view rendering. Note we omit the loss on recorded trajectory [35] in the whole context for clarity.
Algorithm 1 Iterative refinement
0:  Initial Gaussian model $\mathcal{G}_{0}$, novel trajectories $\mathcal{T}^{\prime}$, video diffusion $\mathcal{D}$, total steps $T$, warm up steps $T_{0}$
1:  for  $t=0,\cdots,T-1$ do
2:     $\mathcal{V}_{t}\leftarrow\left\{R(\mathcal{G}_{t},P_{i})\right\}_{i=0}^{F}$
3:     Computing loss $\mathcal{L}_{\text{img}}(\mathcal{V}_{t},\mathcal{V}_{\text{gt}})$
4:     if $t\geq T_{0}$ then
5:        $\mathcal{V}^{\prime}_{t}\leftarrow\left\{R(\mathcal{G}_{t},P_{i}^{\prime})\right\}_{i=0}^{F}$
6:        if $(t-T_{0})\mod K=0$ then
7:           $k\leftarrow t$
8:           $\mathcal{V}^{\prime}_{k,\mathrm{refine}}\leftarrow\mathcal{D}\left(\mathcal{V}^{\prime}_{t},\mathcal{M}\right)$
9:        end if
10:        Computing loss $\mathcal{L}^{\prime}\leftarrow\mathcal{L}_{\text{img}}(\mathcal{V}^{\prime}_{t},\mathcal{V}^{\prime}_{k,\mathrm{refine}})$
11:     end if
12:     Backwarding loss and updating $\mathcal{G}_{t+1}$
13:  end for
14:  return  $\mathcal{G}_{T}$
To assess geometry accuracy in novel view rendered images, we derive $\mathcal{M}$ by comparing the rendered images with novel view warped images from the recorded trajectory. Let $I_{\text{ren}}\in\mathcal{V}^{\prime}_{t}$ denote a rendered image, and let $I_{\text{rec}}$ be the recorded image that are closest to $I_{\text{ren}}$. We define a warping operation $\psi$ in Eq. 7 to project a 3D point $p$ onto the image plane given camera pose $P$:
$(x,y,d)=\psi(p|P),$ (7)
where $x,y$ represent the pixel coordinates, and $d$ is the depth (Please refer to Appendix for details). Notably, $\psi$ is invertible, allowing us to unproject an image with its corresponding depth map back to 3D points given the camera pose. Consequently, we can obtain a pseudo image $\hat{I}_{\text{ren}}$ under the pose $P_{\text{ren}}$ by sampling the color in $I_{\text{rec}}$ using the warped pixel coordinates $(\mathbf{x},\mathbf{y},\mathbf{d})$ from unwarped $I_{\text{ren}}$ to $I_{\text{rec}}$, as expressed by Eq. 8.
$(\mathbf{x},\mathbf{y},\mathbf{d})=\psi(\psi^{-1}(I_{\text{ren}},D_{\text{ren}}|P_{\text{ren}})|P_{\text{rec}}),$ (8)
where $D_{\text{ren}}$ is the rendered depth corresponding to $\hat{I}_{\text{ren}}$. Since it is hard to guarantee the pixel-wise correspondence between $I_{\text{ren}}$ and $\hat{I}_{\text{ren}}$, we choose SSIM to evaluate similarity at the patch-structure level. The unreliability mask $\mathcal{M}$ is then obtained by applying a threshold $\tau$ to SSIM score:
$\mathcal{M}=\mathbbm{1}\left(\text{SSIM}(I_{\text{ren}},\hat{I}_{\text{ren}})<\tau\right).$ (9)
Because the warping process involves both rendered depth and rendered image, the discrepancy can serve as an indicator of geometric or appearance unreliability.
Since the warped image provides a rough overview about the novel trajectory, making it unnecessary to start denoising from scratch when we refine the videos using Eq. 5. Hence, we perturb $\hat{I}_{t}$ by adding Gaussian noise at some noise level. This consideration not only reduces unnecessary time costs but also avoids color shifts by preserving accurate low-frequency components in original video.
EmerNeRF [36] PVG [4] StreetGaussian [35] DriveX (Ours)
Recorded image DriveDreamer4D [40] DriveX (Ours)
Figure 3: Qualitative comparisons of novel trajectories on Waymo. Top: For each method, images are rendered to the left of the recorded trajectory by $3m$. Bottom: Our results are consistent with the recorded image compared to DriveDreamer4D [40].
Iterative refinement
We have detailed how to use generative prior to solve a video inverse problem to improve the reconstruction quality. In turn, the generation can also benefit from these improvements, as it is conditioned on $\mathcal{G}_{t}$ whose enhancements could provide more accurate information for generation. This reciprocal enhancement motivates an iterative refinement. As illustrated in Algorithm 1, we first warm up the optimization for $T_{0}$ steps using conventional reconstruction method [35] on recorded trajectory, and then store the refined video $\mathcal{V}_{t,\text{refine}}^{\prime}$ as a buffer, which will be updated every $K$ steps. Due to the additional time demands of running video diffusion model, it is inefficient to set $K=1$, i.e., the buffer is updated every step. Please refer to Sec. 4.3 for the trade-off between different $K$. Since the generated contents are conditioned on Gaussian model $\mathcal{G}_{t}$, they are constrained with the recorded scene, thus ensuring the consistency between reconstruction and generation.
Novel trajectory sampling
Another important design space is how to sample the novel trajectory $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$. The ideal trajectory should balance the following two requirements: first, maximizing the utility of the generative model, i.e., the generated novel views should contribute as much as possible to enhance reconstruction quality; second, minimizing inconsistency between the recorded and generated images. Therefore, we adopt a panning camera trajectory that starts from a recorded front view $P^{\prime}_{0}=[R_{0}|T_{0}]$ and gradually shifts laterally:
$P^{\prime}_{i}=[R_{0}|\frac{i}{F}s\mathbf{v}+T_{0}],$ (10)
where $\mathbf{v}\in\mathbb{R}^{3}$ is the shifting direction, and $s\in\mathbb{R}$ controls the maximum shifting length. This design not only provides novel views beyond the recorded trajectory, but also allows video generation model to extract detailed references from the initial frame. At the early stage of optimization, rendering quality may deteriorate as the camera diverges from the recorded trajectory, which offers limited guidance for video generation. To address this issue, we initially use a small shifting length $s$ and progressively extend it during optimization, ensuring our method’s stability and robustness across diverse scenes.
±$\pm 1m$$\pm 2m$$\pm 3m$ Inference speed
[…]
EmerNeRF [36] 0.1676 0.5149 91.40 0.1372 0.4671 111.67 0.1369 0.4198 132.48 0.12
$S^{3}$aussian [12] 0.1423 0.5301 70.71 0.0865 0.4781 100.29 0.0430 0.4341 126.22 8.4
PVG [4] 0.1335 0.5406 71.63 0.0632 0.4686 102.92 0.0365 0.4267 137.12 48
StreetGaussian [35] 0.2002 0.5945 49.44 0.1668 0.5793 80.05 0.1385 0.5342 105.37 34
DriveX (Ours) 0.2065 0.5949 48.65 0.1892 0.5827 76.44 0.1755 0.5427 95.60 34
Table 1: Novel trajectory evaluation on Waymo. We compare images rendered in novel trajectories with different shift lengths, where the IoU is computed between the ground truth and those detected by TwinLiteNet [3]. Once trained, our model solely relies on the Gaussians for inference. The inference speeds of all methods are measured on the NVIDIA RTX A6000 GPU.
Gaussian model Lane change Acceleration Deceleration Average
[…]
DriveDreamer4D [40] PVG [4] 0.428 53.00 0.411 53.10 0.421 53.78 0.420 53.29
DriveX (Ours) PVG [4] 0.474 58.53 0.563 65.36 0.611 66.97 0.549 63.62
DriveX (Ours) StreetGauss. [35] 0.516 62.77 0.569 65.14 0.599 66.23 0.561 64.71
Table 2: Comparision with DriveDreamer4D [40] under its evaluation setting. For fair comparison, we report the same metric proposed in [40] measured on the same sequences, and same Gaussian model [4] with the best reported results in [40].
3.3 Application on generated videos
In addition to real-world scenes, the driving videos used for reconstruction can also be created by a video generator [7], which explores new possibilities for the automatic creation of diverse virtual driving environments. However, the AI-generated videos present more severe challenges because of the lack of corresponding LiDAR depth and accurate pixel-to-pixel matching between different frames. These challenges further necessitate the novel view supervision provided by the generative priors.
Therefore, we devise an additional recipe for the robust optimization of the AI-generated videos. Specifically, since the generated videos cannot be precisely aligned with the conditioned camera parameters, we employ a feed-forward approach [39] to estimate the camera parameters and corresponding depths for each frame. For scenes containing moving vehicles, we manually annotate their tracklets as the initialization for the local coordinate frames. During optimization, we fine-tune the camera pose to mitigate potential prediction errors. In the computation of unreliability, as the generated video is monocular, we use the nearest previous frames that can include the current shifted view content as the auxiliary source images in Eq. 8 to ensure well-defined unreliability scores for all areas of the novel view.
4 Experiments
4.1 Experimental setup
Datasets
For evaluation of the improvement for real-world reconstruction, we conduct experiments on 12 selected sequences comprising surrounding videos and synchronized LiDAR point clouds from Waymo Open Dataset [26]. Official tracklets are used to crop the point cloud of dynamic traffic participants and initialize their per-frame pose. For fair comparison, all images are downsampled into $640\times 920$ in both training and testing. To test the applicability of DriveX in generated driving videos, we assess our method on 12 driving scene videos produced by video diffusion models [7]. Each video is generated given the initial frame from nuScenes validation set as the condition.
Metrics
To reasonably benchmark the improvement on the existing reconstruction-based street view synthesis pipeline, we quantitatively compared DriveX with several representative baselines on views from the novel trajectory. Since the ground truth sensor data on these trajectories is unavailable, it is infeasible to directly assess the pixel-level fidelity. Therefore, we devised a novel benchmark to comprehensively evaluate the synthesis results at these views. Specifically, the Fréchet Inception Distance (FID) [9] between synthesized views in novel trajectories and the captured images from the original trajectories is employed to assess the realism of synthesized images at the distribution level. Moreover, we report Lane IoU and vehicle AP to evaluate the distinguishability and fidelity of two safety-critical high-level traffic components, i.e., road lanes and vehicles, in novel trajectory viewpoints. Refer to the appendix for detailed calculations of these metrics.
Implementation details
For reconstruction of real video where calibrated data and LiDAR are available, we first warm up the optimization with Gaussian model [35] for 50,000 steps and then integrate generative priors for following 30,000 steps. For novel trajectory videos, the initial view $P_{0}^{\prime}$ of a novel trajectory $\mathcal{T}^{\prime}$ is selected from every 3 recorded forward cameras, with shifting direction $\mathbf{v}$ oriented to the left or right, and shifting length $s$ ranging from $2m$ to $4m$. During the iterative refinement, the buffer for storing refined videos is updated every $K=2000$ steps. We use ViewCrafter [37] as our generative prior. The refine strength (i.e. the level of noise adding to images) is set to 0.6. The threshold $\tau$ in Eq. 9 is set to 0.65. To alleviate the conflicts between recorded data and generated results on moving cars, their supervision strength is reduced to 0.2. For generated videos, we first estimate the camera pose [39] and predict the depth and then perform similar process.
4.2 Comparison with state of the art
Results on real-world videos
For comparisons on real-world reconstruction, we primarily focused on improvements in substantial novel views beyond the recorded trajectory. To this end, we synthesized novel driving videos along trajectories shifting from the recorded trajectories by ±$\pm 1m$, ±$\pm 2m$ and ±$\pm 3m$. Table 1 demonstrate our significant improvement on FID, AP and IoU, surpassing the leading StreetGaussian [35] with a 9.3% decrease in FID, 1.6% increase in AP and 21.1% increase in IoU under the largest $3m$ shifting length. These enhancements are visually corroborated in Fig. 3, where safety-critical scene elements can be well synthesized in novel views by our methods. It can be observed that by integrating generative priors, our approach demonstrates robust rendering quality under large trajectory deviations, resulting in more realistic driving environments. Notably, once trained, our model solely relies on the Gaussians for inference, enabling efficient rendering speed.
For fair comparison with DriveDreamer4D [40] which is not open-source, we adopt the same experiment setting (sequences, metric and its best reported Gaussian counterpart [4]). The results are presented in Tab. 2 and bottom of Fig. 3. Our model markedly outperforms DriveDreamer4D with variety of Gaussian models on all metrics, visual quality, and the consistency of traffic components.
w/o ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/ ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Reconstruction results on generated videos. All ground truth videos used for reconstruction are generated by Vista [7]. The first and the second rows show the novel trajectory synthesis results by the baseline [35] and our method, respectively.
Method FID $\downarrow$
±$\pm 0m$$\pm 1m$$\pm 2m$
Recon-only [35] 45.45 96.76 146.61
DriveX (Ours) 45.44 92.77 142.73
Table 3: Quantitative comparison on generated video. We compared the FID score on videos generated by Vista [7].
Results on generated videos
Another intriguing application of the proposed method is transforming AI-generated video into a re-enactable driving world. To quantitatively evaluate this capability, we deploy DriveX and the reconstruction-only baseline [35] on 12 driving videos generated by [7]. The FID of videos rendered across different novel trajectories is reported in Tab. 9. The results indicate that the realism of rendered videos is well preserved within a certain range of shifted length. Compared to the reconstruction-only baseline, our DriveX achieves 4.1% and 2.6% FID reduction in shifting length of ±$\pm 1m$ and ±$\pm 2m$ respectively. To obtain real world distance, we align the predicted depth and LiDAR depth of the first frame by the least square estimation. These advancements are also clearly illustrated in Fig. 4. As can be seen, the baseline method exhibits significant degradation in views that are far from the recorded trajectory.
4.3 Ablation study
[…]
PVG [4] 0.0200 0.5401 118.41
PVG [4] 0.0819 0.6188 101.04
StreetGaussian [35] 0.1217 0.6124 103.08
StreetGaussian [35] 0.2092 0.6257 90.30
Table 4: Ablation study on different Gaussian model. Unless stated otherwise, all ablation experiments are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
[…]
mask all 0.0899 0.5781 107.58
w/o mask 0.1106 0.5965 90.49
w/ mask 0.2092 0.6257 90.30
Table 5: Effects of unreliability mask.
[…]
500 0.1070 0.5954 86.56 4.8×4.8$\times$4.8 ×
2,000 0.2092 0.6257 90.30 1.7×1.7$\times$1.7 ×
5,000 0.1104 0.5992 105.25 1.4×1.4$\times$1.4 ×
Inf 0.0945 0.5845 95.08 1.0×1.0$\times$1.0 ×
Table 6: Ablations on buffer intervals. Inf denotes the novel trajectory supervision only generated once time at iteration $T_{0}$.
(a) Ablation on unreliability mask (b) Ablation on iterative refinement
Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement. Fig.(a) shows the refined results without (middle left) and with (middle right) unreliability mask (top right). The final rendered results exhibit significant artifacts without mask (bottom left). Fig.(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right). It can be seen that single-time refinement can not completely eliminate artifacts, resulting in inferior final rendered results compared to iterative refinement (bottom). $T_{0}$ and $T$ indicate the first and last iteration after warm up, respectively.
Different Gaussian model
Although we adopt StreetGaussian [35] as the default Gaussian counterpart in previous experiments, our method is also capable of synergy with different choices. To demonstrate the versatility of DriveX, we evaluate on a dynamic counterpart PVG [4]. From Tab. 4, we can observe that our strategy consistently reduces the FID by a large margin, and effectively recovering street lanes and cars (i.e. IoU, AP) that are entirely indiscernible in the baseline methods. Unless stated otherwise, all ablation experiments are conducted on two Waymo sequences.
Unreliability mask
The unreliability mask is designed to guide the generation of video diffusion. From Tab. 5 and Fig. 5(a) we can observe: (i) When all regions are masked, where video diffusion model does not know any information from the in-training Gaussian, leading to inferior results; (ii) When the mask is disabled, the generative model can still refine the rendered results toward the video distribution, but with minor refinement; (iii) Equipped with proposed unreliability mask, our full model achieves superior performance in both quantitative and qualitative ablation.
Iterative refinement
The interval of iterative refinement offers the trade-off between quality and efficiency. When the novel trajectory supervision only generated at iteration $T_{0}$ (see Fig. 5(b)), the quality of the novel view is constrained to the in-training Gaussian model which is not fully optimized. To the other extreme, we are unable to afford the time cost of running the diffusion at every step. From Tab. 6, a small interval 500 leads to 4.8×4.8$\times$4.8 × training time. Hence we finally chose the interval 2,000 to strike the balance.
5 Conclusion
In this paper, we present a novel framework that integrates generative priors into the reconstruction of the driving scene from single-trajectory recorded videos, addressing limitations in extrapolating novel views beyond a recorded trajectory. To reasonably leverage the video generative model to provide consistent novel view supervision, we construct a tailored inverse problem to refine the novel view rendered images and progressively enhance the street scene model. The extensive experiments demonstrate that the proposed method significantly improves the synthesis quality for novel trajectories, both in real and AI-generated scenes. These results pave the new way for driving scene synthesis on free-form trajectories.
Supplementary Material
6 More implementation details
Used sequences
The results in Table 1 of the main text are computed from 12 sequences curated from WOD [26]. The names of all sequences are listed below:
  • segment-10444454289801298640
  • segment-10625026498155904401
  • segment-6242822583398487496
  • segment-8822503619482926605
  • segment-10359308928573410754
  • segment-11450298750351730790
  • segment-12496433400137459534
  • segment-15021599536622641101
  • segment-16767575238225610271
  • segment-17860546506509760757
  • segment-3015436519694987712
  • segment-6637600600814023975
For the reconstruction from generated videos, the videos are generated given the following images from nuScenes [2] validation set:
  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537853021662460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984433512460
  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537852899162460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984912912460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984834412460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984833912460
  • n015-2018-07-18-11-41-49+0800__CAM_FRONT__1531885347362460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984560512460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984563012460
  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657697012404
  • n008-2018-09-18-14-35-12-0400__CAM_FRONT__1537295844362404
  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657694012404
Metric details
Previous driving scene reconstruction methods [38, 8, 13] typically utilize PSNR, LIPIS and SSIM [31] for quality evaluation. However, these metrics are inapplicable for evaluating free-form trajectories due to the absence of corresponding ground-truth images for novel trajectories. Latest research [40] proposed Novel Trajectory Agent IoU (NTA-IoU) and Novel Trajectory Lane IoU (NTL-IoU) for traffic components evaluation. Accordingly, we compared NTA-IoU and NTL-IoU, as shown in Table 2.
Furthermore, we propose improvements to the evaluation metrics of NTA-IoU and NTL-IoU. For vehicle quality, similar to NTA-IoU, we projected the original 3D vehicle bounding boxes onto the new viewpoints to generate corresponding 2D bounding boxes. IoU lacks the precision granularity needed to understand performance across various intersection over union threshold. Therefore, we calculate the Average Precision (AP[50:95]) which provides a more nuanced assessment of detection accuracy.
For lane quality, NTL-IoU calculates the mIoU between the rendered and the ground truth lanes, averaging both foreground lanes and backgrounds. However, including the background can inflate accuracy scores due to the easy detection of non-lane areas. Therefore, following official lane detection benchmarks [3] we focus specifically on IoU with only the foreground to better evaluate lane recognition performance.
For the evaluation of the capability as driving the scene generator, we mainly assess the realism of synthesized images across different trajectories by measuring their FID with the images at recorded trajectories.
More details
For reconstruction from real data, we adopt regularization loss as introduced in [35], including sky loss, LiDAR depth loss and entropy loss for objects. In our framework, the LiDAR depth loss is used for both recorded trajectory and novel trajectory. In addition, we balance the recorded and generated data with 50% of the training data being generated. For reconstruction from generated videos, we additionally include an entropy loss for the opacity of each Gaussian to prevent potential overfitting.
Warping operation $\phi$ in Eq. 7
Given a camera extrinsic matrix $E\in\mathbb{R}^{4\times 4}$ and intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, the mapping from world points $p_{\text{world}}\in\mathbb{R}^{3}$ to pixel coordinates $(x,y)$ and depth $d$ can be expressed by:
$\begin{bmatrix}p_{\text{camera}}\\ 1\end{bmatrix}$ $=(*,*,d,1)^{\top}=E\begin{bmatrix}p_{\text{world}}\\ 1\end{bmatrix}$ (11)
$(x_{*},y_{*},d)^{\top}$ $=Kp_{\text{camera}}$ (12)
$(x,y)$ $=(\frac{x_{*}}{d},\frac{y_{*}}{d})$ (13)
Note that all the calculation above is invertible, allowing the mapping from $(x,y,d)$ to world points $p_{\text{world}}$.
7 More ablations
Refine strength
[…]
0.4 0.1122 0.5904 98.31 1.0×1.0$\times$1.0 ×
0.6 0.2092 0.6257 90.30 $\times1$
0.8 0.0996 0.5874 92.26 1.2×1.2$\times$1.2 ×
1.0 0.0773 0.5979 117.91 1.3×1.3$\times$1.3 ×
Table 7: Effect of the noise level. In this table, we evaluate the performance under different refine strengths. The metrics are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
When we refine the rendered videos using video diffusion, we also study the effect the refine strength (level of noise added to the images). In Tab. 7, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time. With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement. On the other hand, a high refine strength yields better novel views but may not keep fidelity to the original scene. Additionally, higher refine strength requires more denoising steps, increasing the training time. Consequently, a refine strength of 0.6 is found to be the best balance.
Other designs
[…]
w/o pose refinement N/A N/A 116.86
w/o depth prior N/A N/A 98.62
Full model N/A N/A 95.73
w/o progress. 0.1075 0.5959 91.22
w/o warp 0.0652 0.5962 119.24
Full model 0.2092 0.6257 90.30
Table 8: Ablations on other design choices. The metric in the first three lines is measured on the generated video, the others are tested on real scene.
We also ablate other design choices in Tab. 8, including the pose refinement and depth prior in the reconstruction from generative video, as well as progressive increasing shifted length and warped condition. The results show that the final quality degrades without any element. The results indicate that the exclusion of any element leads to a decline in the rendering quality and the stability of optimization.
8 Comparison with video generative model
FID $\downarrow$
±$\pm 0m$$\pm 1m$$\pm 2m$
MagicDrive-t [6] 125.07 130.23 138.31
Vista [7] 97.48 99.19 148.07
Recon-only [35] 41.55 95.32 144.01
Ours 36.54 71.47 119.11
Table 9: Comparison with the video-based driving world model. The FID score is calculated between the recorded videos from the nuScenes dataset [2] and the novel trajectory videos synthesized by each method.
In this section, we compare the novel trajectory video synthesis with representative open-sourced generative-based driving world models [7, 6]. Since these works are primarily trained on nuScenes dataset [2], we still conduct the experiments on 12 sequences from this dataset mentioned in Sec. 6. Different from the experiment in Tab. 3, in this experiment, our method takes ground truth recorded trajectory videos as input. The FID between the recorded videos and the synthesized novel trajectory videos from the different methods are reported in Tab. 9.
Note that it is difficult to precisely control the trajectory for generative model. Though we adjust the action control, it still tend to generate videos in the original trajectory. Although all methods do not have significant artifacts, DriveX still demonstrates a clear advantage in FID metrics. This is because it can effectively avoid hallucinations that may be produced by generative models, making its result better aligned with the ground truth. Another advantage difficult to demonstrate by this metric is that DriveX can achieve precise camera pose control. In contrast, other methods are constrained by the diversity and imbalance of actions existing in public driving datasets, thus making it difficult to achieve accurate free-form trajectory synthesis.
9 More results
Figure 6: More comparison results.
Fig. 6 shows more comparison results. We recommend readers to watch the video in the supplementary material.

### v2

@@
-Driving Scene Synthesis on Free-form Trajectories with Generative Prior
+DriveX: Driving View Synthesis on Free-form Trajectories with Generative Prior
 Abstract
-Driving scene synthesis along free-form trajectories is essential for driving simulations to enable closed-loop evaluation of end-to-end driving policies.
-While existing methods excel at novel view synthesis on recorded trajectories, they face challenges with novel trajectories due to limited views of driving videos and the vastness of driving environments.
-To tackle this challenge, we propose a novel free-form driving view synthesis approach, dubbed DriveX, by leveraging video generative prior to optimize a 3D model across a variety of trajectories.
-Concretely, we crafted an inverse problem that enables a video diffusion model to be utilized as a prior for many-trajectory optimization of a parametric 3D model (e.g., Gaussian splatting).
-To seamlessly use the generative prior, we iteratively conduct this process during optimization.
-Our resulting model can produce high-fidelity virtual driving environments outside the recorded trajectory, enabling free-form trajectory driving simulation.
-Beyond real driving scenes, DriveX can also be utilized to simulate virtual driving worlds from AI-generated videos.
-Figure 1: Though conventional reconstruction method can well fit views on the recorded trajectory, they struggle to extrapolate onto the views along the novel trajectories (top).
-In contrast, our method significantly improves the synthesis quality on these novel views (bottom), thereby achieving driving scene synthesis on free-form trajectories.
+Driving view synthesis along free-form trajectories is essential for realistic driving simulations, enabling closed-loop evaluation of end-to-end driving policies.
+Existing methods excel at view interpolation along recorded paths but struggle to generalize to novel trajectories due to limited viewpoints in driving videos.
+To tackle this challenge, we propose DriveX, a novel free-form driving view synthesis framework, that progressively distills generative prior into the 3D Gaussian model during its optimization.
+Within this framework, we utilize a video diffusion model to refine the degraded novel trajectory renderings from the in-training Gaussian model, while the restored videos in turn serve as additional supervision for optimizing the 3D Gaussian.
+Concretely, we craft an inpainting-based video restoration task, which can disentangle the identification of degraded regions from the generative capability of the diffusion model and remove the need of simulating specific degraded pattern in the training of the diffusion model.
+To further enhance the consistency and fidelity of generated contents, the pseudo ground truth is progressively updated with gradually improved novel trajectory rendering, allowing both components to co-adapt and reinforce each other while minimizing the disruption on the optimization.
+By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time—unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
+Figure 1: Though conventional reconstruction-based approaches can well fit views on the recorded trajectory, they struggle to extrapolate onto the views beyond it (top).
+In contrast, our method significantly improves the synthesis quality on these novel trajectory views by incorporating generative prior (bottom), thereby achieving superior synthesis quality on free-form trajectories.
 1 Introduction
-Building virtual driving worlds engagible with driving policies plays a pivotal role in developing robust autonomous driving systems.
-It enables automatic synthesis of diverse driving data, including safety-critical long-tail scenarios, edge cases, and providing closed-loop evaluation for end-to-end autonomous driving systems.
-Recent advancements in novel view synthesis [21, 15] have driven reconstruction-based driving simulation [34, 4, 35] due to its ability to synthesize photorealistic sensor data.
-These methods aim to reconstruct driving environments from individual driving sequences, typically involving single-trajectory videos from surrounding cameras with sparse views, minimal overlap, and textureless areas.
-Despite impressive results in fitting training views and interpolating along original trajectories, they still struggle to extrapolate substantially novel views far from the recorded trajectory, leading to limited flexibility and usefulness for driving simulation.
-To address these problems, a couple of latest works [13, 38] explored image generative prior as extra novel view regularization.
-However, they are largely limited in design due to (i) the need for fine-tuning an off-the-shelf image generation models to obtain the capability of view transformation, suffering from both additional computational cost and per-scene data scarcity, and (ii) absence of the desired spatiotemporal knowledge within image models.
-To tackle these issues, a concurrent work [40] uses a pretrained video diffusion model to generate novel trajectory training images conditioned on initial recorded frames.
-Without sufficient content constraint but with intrinsic hallucination risk, this approach however tends to drift away from the original scene in scene generation, making troubles for the subsequent model optimization.
-To address the challenges mentioned above, we propose a novel driving scene synthesis approach for free-form trajectories, termed DriveX, by leveraging video generative prior with rich spatiotemporal knowledge during the model (e.g., Gaussian splatting) optimization.
-This is achieved by designing an inverse problem that enables a video diffusion model to serve as a prior for optimizing the Gaussian over a variety of trajectories.
-During training, we render novel trajectory views using the in-training model.
-To identify potential artifacts from trajectory extrapolation, we compare each rendered view with recorded images wrapped under the same view and geometric conditions, allowing us to obtain reliable regions for each rendered image.
-The inverse problem is then crafted as restoring each rendered view via utilizing its corresponding reliable regions as conditioning inputs, which can be effectively executed by the video diffusion model.
-The output from this process is subsequently fed into the optimization as supervision to update the target model.
-To seamlessly exploit the generative prior, we iteratively conduct this process during Gaussian optimization.
+Building virtual driving worlds capable of being engaged with driving policies plays a pivotal role in developing robust autonomous driving systems.
+It facilitates efficient training data scaling and automatic synthesis of safety-critical long-tail cases, while enabling close-loop evaluation for end-to-end autonomous driving systems.
+In recent years, the burgeoning advancements in novel view synthesis [19, 13] have propelled the prevalence of reconstruction-based driving simulations [36, 3, 37].
+These methods aim to reconstruct driving environments from vehicle sensor data typically involving single-trajectory videos from surrounding cameras and synchronized LiDAR point clouds.
+However, the limited overlap between frames, large textureless regions and illumination variations in large-scale scenes, drastically complicate this task than conventional novel view synthesis.
+Consequently, despite achieving impressive results in fitting training views and interpolating along recorded trajectories, they still struggle to extrapolate on novel views far from the recorded trajectory, limiting their flexibility and usefulness in many downstream tasks.
+Recent works [16, 17, 39] have demonstrated that 2D generative priors can effectively regularize sparse-view reconstruction by synthesizing pseudo ground-truth images for novel view, thereby alleviating the under-constraint nature of this task.
+However, directly applying these approaches in the driving scenes presents unique challenges due to the limited diversity of trajectories in existing driving video datasets.
+The training is primarily confined on forward-moving trajectories, while the off-trajectory videos are required for inference.
+This will introduce a considerable gap, making it infeasible to learn precise explicit pose control, and posing a great challenge to simulate degraded patterns for training a restoration-based model.
+To address these challenges, we propose a reconstruction-generation interwined framework for driving view synthesis along free-form trajectories, termed DriveX, which progressively distills video generative prior into the 3D scene representation (i.e., 3D Gaussian Splatting [13]).
+During this process, a video diffusion model is employed to generate novel trajectory videos as additional supervision for the optimization of scene representation, effectively mitigating the sparse-view issue.
+This is achieved by leveraging a generative model to restore novel trajectory renderings from the in-training 3D Gaussian model, which ensures the consistency of generated content while eliminating the need for explicit pose condition.
+Specifically, we design an inpainting-based restoration task where generation is conditioned on the rendered video masked by a geometry-aware unreliability mask.
+This formulation allows the model to fully leverage its generative capability, freeing it from the task of identifying degraded regions, thus circumventing the complexity of simulating degraded patterns during its training.
+During the optimization of scene representation, we progressively perform the restoration process and update the generated pseudo ground truth to further enhance consistency of the generated content.
 As shown in Fig.
-1, our approach substantially improve the view synthesis quality in novel trajectory compared to the traditional reconstruction methods.
-The contributions of this paper are as follows: (i) We propose leveraging video generative prior with rich spatiotemporal knowledge for generalizable driving scene synthesis from single-trajectory recorded videos.
-(ii) We introduce DriveX, a novel driving scene synthesis framework for free-form trajectories that innovatively constructs an inverse problem to enable the use of a video diffusion model as a prior.
-(iii) Extensive experiments demonstrate that DriveX significantly outperforms existing state-of-the-art alternatives in driving scene synthesis using single-trajectory recorded videos.
-Additionally, we showcase the superiority of our method in rendering novel trajectories of virtual driving worlds from AI-generated videos, even when faced with inherent content inconsistencies.
-This approach enables a more economically scalable large-scale simulation without the need to collect videos for each scene.
+1, our approach substantially improves the view synthesis quality in the novel trajectory compared to the state-of-the-art reconstruction-only method [37].
+Our contributions are as follows: (i) We propose a novel driving view synthesis framework that seamlessly integrates generative prior into driving scene reconstruction, where a diffusion model is employed to progressively synthesize consistent novel view ground truth for regularizing the optimization of the 3D Gaussian model; (ii) By formulating an inpainting-based restoration task, we disentangle the identification of degraded regions from the generative capability of the diffusion model, thereby simplifying training on driving videos with limited diverse trajectories; (iii) Our DriveX effectively overcomes the sparse-view issue in previous street scene reconstruction methods, enabling high-fidelity view synthesis along free-form trajectories.
 2 Related works
-Reconstruction-based driving scene synthesis
-Early approaches [34, 36, 27, 32, 28] for dynamic urban scenes reconstruction focus on extending neural radiance fields (NeRF) [21] in large scale unbounded scenes.These methods are restricted with low efficiency and blurry rendering performance.
+Reconstruction-based driving view synthesis   Early approaches [36, 38, 26, 32, 27] for dynamic urban scenes reconstruction focus on extending neural radiance fields (NeRF) [19] in large scale unbounded scenes.
+These methods are restricted with low efficiency and blurry rendering performance.
 Recently, another line of works introduced 3D Gaussian splatting [13] into this task to produce real-time high-fidelity rendering.
@@
 However, driving scenes featured with sparse views, little overlaps, and large textureless areas, making existing methods struggle to generalize to substantively novel views.
-Generative-based driving scene synthesis
-Recently, propelled by the significant advancement of the diffusion model [10, 24, 22, 25, 1], many works employ video generators to construct world models [16].
+Generative-based driving view synthesis   Recently, propelled by the significant advancement of the diffusion model [9, 23, 22, 24, 1], many works employ video generators to construct world models [14].
 Some of them introduce this paradigm into autonomous driving [29, 45, 15, 5, 30, 6, 10, 18] for simulating the plausible visual outcomes under versatile driving controls.
@@
 Consequently, they cannot promise geometry and texture consistency for the generated many-trajectory videos of the same scene.
-Figure 2: The schematic illustration of DriveX.
-Top: Gaussian optimization with both recorded trajectory and generated novel trajectory as supervision.
+Figure 2: Schematic illustration of DriveX.
+Top: Optimizing the scene representation with both recorded trajectory views and generated novel trajectory views as supervision.
 Bottom: Generative prior integration.
-Given a novel view rendered image from Gaussians $\mathcal{G}_{t}$, video diffusion is applied to generate its refined version with condition of unreliability mask.
-We omit the loss on recorded trajectory for clarity.
-Scene reconstruction with diffusion prior
-Leveraging generative prior to enhance the reconstruction from sparse view observations [37, 19, 18] has emerged.
+During optimization, given a novel view rendering from in-training Gaussian $\mathcal{G}_{t}$, a video diffusion model is employed to generate a restored version as pseudo ground truth on novel trajectories, guided by its unreliability mask and accumulated LiDAR projections.
+Scene reconstruction with diffusion prior   Leveraging generative prior to enhance the reconstruction from sparse view observations [39, 17, 16, 34] has emerged.
 Extending this paradigm into driving scene reconstruction has been recently attempted [40, 12, 44].
-However, they typically suffer from limited diversity and scale of driving scene data.
+Limited diversity of trajectory in driving scene data poses unique challanges for them.
 SGD [40] and VEGS [12] contribute to define and tackle pseudo and extrapolated view synthesis in driving scenarios.
 However, these methods may encounter challenges with accurate pose and detail estimation under novel trajectories.
-GGS [8] introduces virtual lane generation module, but its dependency on modified MVSplat structure poses challenges for generalization to large-scale pretraining.
-A concurrent work [40] uses a world model [29] to generate novel trajectory video referenced on initial frame for trajectory augmentation.
-However, it tends to drift away from the original scene due to the strong built-in hallucination risk and insufficient constraints.
-This would inevitably hurt the subsequent model optimization.
-In contrast, our method dedicates to solve an inverse problem that enables a video diffusion model to be used as a prior for many-trajectory optimization of a parametric 3D model, achieving superior synthesis quality on free-form trajectories.
+DriveDreamer4D [44] samples novel trajectory video based on single reference frame before training, risking hallucination and inconsistency due to insufficient constraints.
+StreetCrafter [41] refines the novel view renderings using a controllable video diffusion model guided by LiDAR projections.
+Meanwhile, FreeSim [4] and Difix3D+ [33] directly condition on degraded renderings and employ carefully data curation to train the generative model to adapt on degradation patterns resulted from off-trajectory renderings.
+In contrast, by crafting an inpainting-based video restoration task, we free the generative model from identifying degraded regions, thereby eliminating the need to simulate specific degraded patterns in training.
 3 Methodology
@@
 Reconstruction-based driving simulation aims to recover the surrounding scenes from calibrated video $\mathcal{V}_{\mathrm{gt}}=\left\{I_{i}\right\}_{i=0}^{F}$ recorded along a driving trajectory $\mathcal{T}=\left\{P_{i}\right\}_{i=0}^{F}$, where $P_{i}\in\mathrm{SE(3)}$ is the camera pose at frame $i$.
-Then the sensor data would be rendered when the vehicle performs simulated unrecorded actions (e.g.
+Then the sensor data can be simulated under the unrecorded actions (e.g.
 lane change) in the reconstructed environment.
@@
 where $\mathcal{L}_{\text{SSIM}}$ is the SSIM [31] loss and $\lambda$ is a weight.
-In the urban scenes, although the traffic participants are dynamic, they can be regarded as static in their own local coordinate frames.
-The transformation from each object’s local frame to the global coordinate frame of the whole scene is parameterized by a series of rotations and transformations, which can be derived from its tracklet annotated by off-the-shelf 3D trackers [5, 33].
-Then a complete dynamic street scene can be modeled by decoupled foreground moving Gaussians and static background Gaussians.
-Additionally, cube map is commonly used to model the sky with infinite distance [35, 4].
-Challenges
-Since the camera configuration in autonomous vehicles is primarily designed for perception tasks, the captured videos often offer limited viewpoints, minimal overlap, and large textureless regions.
+Following [37], we model the traffic participants in the driving scenes as a dynamic node in the neural scene graph with rigid motion, which can be regarded as static in their own local coordinate frames.
+The transformation from each object’s local frame to the global coordinate frame of the whole scene is parameterized by a series of rotations and transformations, which can be annotated by their tracked bounding boxes.
+Additionally, cube map is employed to model the sky with infinite distance [37, 3].
+Challenges   Since the camera configuration in autonomous vehicles is primarily designed for perception tasks, the captured videos often offer limited viewpoints, minimal overlap, and large textureless regions.
 These inherent properties bring challenges to the reconstruction of the unbounded driving scenes from single-trajectory videos.
 Fitting the training views at recorded trajectory $\mathcal{T}$ is insufficient to realize satisfactory extrapolation out of $\mathcal{T}$ for novel view synthesis.
-As a result, existing optimization-based driving scene synthesis approaches struggle to achieve high-quality view rendering along novel trajectories $\mathcal{T}^{\prime}=\{P^{\prime}_{i}|P^{\prime}_{i}\notin\mathcal{T}\}_{i=0}^{F}$, often leading to obvious degeneration and artifacts.
-3.2 DriveX
-To address the aforementioned challenges, we propose a novel framework, DriveX, seamlessly integrating generative priors into the driving scene reconstruction process (e.g.
-Gaussian optimization).
-As shown in Fig.
-2, it iteratively synthesizes novel trajectory scenes by utilizing a video diffusion as supervision to optimize the Gaussian model.
-To ensure the generated content aligns with the underlying scenes in detail, the generation process is appropriately conditioned on the in-training Gaussian model.
-To achieve this, we delicately design an inverse problem [14, 23], which can be formulated as recovering clean novel view images $\mathcal{V}$ from the artifact-heavy novel view rendered images $\mathcal{V}^{\prime}$ on novel trajectories $\mathcal{T}^{\prime}$.
-Concretely, $\mathcal{V}^{\prime}$ can be regarded as a measurement of unknown ground truth $\mathcal{V}$, as:
-$\mathcal{V}^{\prime}=f(\mathcal{V})+\epsilon,$ (3)
-where $\epsilon$ is a random noise, and the measurement function $f$ distorts $\mathcal{V}$ due to poor generalization ability of Gaussian model in conventional methods.
-With this design, a video generative model [37] can be readily employed to address this problem effectively.
-The refined video can then serve as additional supervision within Gaussian optimization.
-Solving inverse problem with generative prior
-The key to solve the inverse problem lies in accurately identifying the artifacts in the novel view rendered video.
-Inspired from the sparse view reconstruction methods [34, 19], we construct an unreliability mask representing the artifacts via comparing the rendered images with recorded images, which is used to directly guide the diffusion model focus on refining these unreliable regions while preserving reliable areas.
-Specifically, given a sequence of viewpoints $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ along a novel trajectory, a video diffusion model $\mathcal{D}$ is employed to estimate $\mathcal{V}_{i}$ in Eq.
-3:
+As a result, existing optimization-based driving scene synthesis approaches struggle to achieve high-quality rendering along novel trajectories $\mathcal{T}^{\prime}=\{P^{\prime}_{i}|P^{\prime}_{i}\notin\mathcal{T}\}_{i=0}^{F}$, often leading to obvious degeneration and artifacts.
+3.2 Novel view regularization with generative prior
+Generate consistent novel view via video restoration A promising solution to the aforementioned challenges is to incorporate generative priors as an additional regularization.
+Large video generative models trained on internet-scale data have developed a strong knowledge of natural video distributions, enabling them to reliably infer and generate novel view content as pseudo ground truth for optimizing 3DGS.
+To ensure consistency between the generated content and the underlying scene, thereby preventing disruption to optimization, the generation should be appropriately conditioned on the in-training Gaussian model.
+To achieve this, we delicately design a video restoration task, which is formulated as recovering clean novel view images $\mathcal{V}$ from the artifact-heavy novel view renderings $\mathcal{V}^{\prime}$ beyond the recorded trajectory.
+Within this design, a video generative model can be employed to address this problem effectively.
+The refined video then serves as additional supervision in the optimization of 3DGS.
+Concretely, given a sequence of viewpoints $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ along a novel trajectory, a video diffusion model $\mathcal{D}$ is employed to estimate clean version $\mathcal{V}_{i}$:
 $\mathcal{V}^{\prime}_{t}$ =$=\left\{R(\mathcal{G}_{t},~{}P_{i}^{\prime})\right\}_{i=0}^{F},$ (3)
-$\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ $=\mathcal{D}\left(\mathcal{V}^{\prime}_{t},~{}\mathcal{M}\right),$ (5)
+$\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ $=\mathcal{D}\left(\mathcal{V}^{\prime}_{t}\right),$ (4)
 $\mathcal{L}^{\prime}$ $\triangleq\mathcal{L}_{\text{img}}(\mathcal{V}^{\prime}_{t},~{}\mathcal{V}^{\prime}_{t,\mathrm{refine}}),$ (5)
-where $\mathcal{V}^{\prime}_{t}$ is rendered from the Gaussian model $\mathcal{G}_{t}$ at current iteration $t$ through Eq.
-1, the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ can be regard the estimation of $\mathcal{V}$ in Eq.
-3, and $\mathcal{M}$ is the mask indicating unreliable regions.
-Typically, $\mathcal{D}$ takes noise-perturbed images as input, conditioned on the masked video, and outputs the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$.
-Then $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ is used as the additional supervision for the novel trajectory loss $\mathcal{L}^{\prime}$ (Eq.
-6) to improve the novel view rendering.
-Note we omit the loss on recorded trajectory [35] in the whole context for clarity.
+where $\mathcal{V}^{\prime}_{t}$ is rendered from the Gaussian model $\mathcal{G}_{t}$ at $t$-th iteration of optimization through Eq.
+1, the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ can be regarded as the estimation of $\mathcal{V}$.
+Typically, $\mathcal{D}$ takes noise-perturbed images as input, conditioned on the degraded video, and outputs the refined video $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$.
+Then $\mathcal{V}^{\prime}_{t,\mathrm{refine}}$ is used as the pseudo ground truth in the novel trajectory loss $\mathcal{L}^{\prime}$ (Eq.
+5) to supervise the novel view renderings as illustrated in Fig.
+2.
 Algorithm 1 Iterative refinement
@@
 14:  return  $\mathcal{G}_{T}$
-To assess geometry accuracy in novel view rendered images, we derive $\mathcal{M}$ by comparing the rendered images with novel view warped images from the recorded trajectory.
-Let $I_{\text{ren}}\in\mathcal{V}^{\prime}_{t}$ denote a rendered image, and let $I_{\text{rec}}$ be the recorded image that are closest to $I_{\text{ren}}$.
+Iterative refinement The generative model can be employed to provide novel view supervision for enhancing the novel view renderings.
+In turn, the generation quality also benefits from these improvements, as it is conditioned on $\mathcal{G}_{t}$ whose enhancements could provide more accurate information to guide the generation process.
+This co-adapt and reinforced enhancement motivates an iterative refinement in our framework.
+As outlined in Algorithm 1, we initiate the optimization for $T_{0}$ steps using conventional reconstruction-based method [37].
+Subsequently, we perform video restoration in Eq.
+4 based on the in-training Gaussians $\mathcal{G}_{T_{0}}$, and use a buffer to store refined videos for reuse in the following $K$ iterations.
+To ensure that the generated content can be improved along with the progressively optimized reconstruction results, we update all videos in the buffer every $K$ iterations by performing video restoration conditioned on the latest novel trajectory renderings.
+Due to the additional time demands of running the video diffusion model, it is inefficient to set $K=1$, i.e., the buffer is updated in every iteration.
+We set it to $3000$ to strike the balance between quality and efficiency.
+Please refer to Sec.
+4.4 for the trade-off between different choice of $K$.
+As the generated contents are conditioned on Gaussian model $\mathcal{G}_{t}$, they are constrained with the recorded scene, thus ensuring the consistency between the recorded and generated ground truth and leading to stable optimization.
+Novel trajectory sampling Another important design space is how to sample the novel trajectory $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$ to be distilled by video diffusion model.
+The ideal trajectory should balance the two requirements: (i) maximizing the utility of the generative model, i.e., the generated novel views should contribute as much as possible to enhance reconstruction quality; (ii) minimizing inconsistency between the recorded and generated images to mitigate potential perturbation for Gausssian’s optimization.
+Therefore, we adopt a panning camera trajectory that starts from a recorded front view $P^{\prime}_{0}=[R_{0}|T_{0}]$ and gradually shifts laterally:
+$P^{\prime}_{i}=[R_{0}|\frac{i}{F}s\mathbf{v}+T_{0}],$ (6)
+where $\mathbf{v}\in\mathbb{R}^{3}$ is the shifting direction, and $s\in\mathbb{R}$ controls the maximum shifting length.
+This design not only provides novel views beyond the recorded trajectory, but also allows video generation model to extract detailed references from the initial frame.
+At the early stage of optimization, rendering quality may degrade as the camera deviates from the recorded driving trajectory, which offers limited guidance for the video generation.
+To address this issue, we initially use a small shifting length $s$ and progressively extend it during optimization, ensuring our method’s stability and robustness across diverse scenes.
+3.3 Inpainting-based video restoration
+Under the proposed trajectory sampling scheme with progressively increasing shifted length, degradations in the rendered video caused by the poor extrapolation capabilities of the scene representation are typically localized.
+Therefore, the key to effective video restoration lies in accurately identifying the artifacts in the rendered videos.
+Ideally, a well-trained generative model should be capable of automatically identifying and restoring these degraded regions.
+However, limited viewpoints in driving video datasets make it challenging to simulate degradation caused by off-trajectory rendering while obtaining corresponding ground truth during the training of the generative model.
+Therefore, we propose an inpainting-based restoration framework that removes the need for the model to adapt to specific degraded patterns, thereby simplifying the training process.
+Specifically, we construct an unreliability mask representing the artifacts via comparing the rendered images with the recorded images, which directly guides the model to focus on refining these unreliable regions while preserving reliable areas.
+Thus, the Eq.
+4 should be revised as:
+$\mathcal{V}^{\prime}_{t,\mathrm{refine}}=\mathcal{D}\left(\mathcal{V}^{\prime}_{t},~{}\mathcal{M}\right),$ (7)
+where $\mathcal{M}$ is the mask indicating unreliable regions, and $\mathcal{D}$ is conditioned on the masked video.
+Unreliability mask To assess geometry accuracy in novel view rendered images, we derive $\mathcal{M}$ by comparing the rendered images with novel view warped images from the recorded trajectory.
+Let $I_{\text{ren}}\in\mathcal{V}^{\prime}_{t}$ denote a rendered image, and let $I_{\text{rec}}$ represent the recorded image that are “closest” to $I_{\text{ren}}$.
 We define a warping operation $\psi$ in Eq.
@@
 $(x,y,d)=\psi(p|P),$ (8)
-where $x,y$ represent the pixel coordinates, and $d$ is the depth (Please refer to Appendix for details).
+where $x,y$ represent the pixel coordinates, and $d$ is the depth (Please refer to Appendix B for details).
+Note that the refined video starts from a front view on the recorded trajectory, which is selected as the “closest” image for all refined views in the video.
 Notably, $\psi$ is invertible, allowing us to unproject an image with its corresponding depth map back to 3D points given the camera pose.
 Consequently, we can obtain a pseudo image $\hat{I}_{\text{ren}}$ under the pose $P_{\text{ren}}$ by sampling the color in $I_{\text{rec}}$ using the warped pixel coordinates $(\mathbf{x},\mathbf{y},\mathbf{d})$ from unwarped $I_{\text{ren}}$ to $I_{\text{rec}}$, as expressed by Eq.
-8.
-$(\mathbf{x},\mathbf{y},\mathbf{d})=\psi(\psi^{-1}(I_{\text{ren}},D_{\text{ren}}|P_{\text{ren}})|P_{\text{rec}}),$ (8)
-where $D_{\text{ren}}$ is the rendered depth corresponding to $\hat{I}_{\text{ren}}$.
+9.
+$(\mathbf{x},\mathbf{y},\mathbf{d})=\psi(\psi^{-1}(D_{\text{ren}}|P_{\text{ren}})|P_{\text{rec}}),$ (9)
+where $D_{\text{ren}}$ is the depth map corresponding to $\hat{I}_{\text{ren}}$ rendered by $\alpha$-blending each Gaussian center’s depth.
 Since it is hard to guarantee the pixel-wise correspondence between $I_{\text{ren}}$ and $\hat{I}_{\text{ren}}$, we choose SSIM to evaluate similarity at the patch-structure level.
@@
 Because the warping process involves both rendered depth and rendered image, the discrepancy can serve as an indicator of geometric or appearance unreliability.
-Since the warped image provides a rough overview about the novel trajectory, making it unnecessary to start denoising from scratch when we refine the videos using Eq.
-5.
-Hence, we perturb $\hat{I}_{t}$ by adding Gaussian noise at some noise level.
-This consideration not only reduces unnecessary time costs but also avoids color shifts by preserving accurate low-frequency components in original video.
-EmerNeRF [36] PVG [4] StreetGaussian [35] DriveX (Ours)
-Recorded image DriveDreamer4D [40] DriveX (Ours)
-Figure 3: Qualitative comparisons of novel trajectories on Waymo.
-Top: For each method, images are rendered to the left of the recorded trajectory by $3m$.
-Bottom: Our results are consistent with the recorded image compared to DriveDreamer4D [40].
-Iterative refinement
-We have detailed how to use generative prior to solve a video inverse problem to improve the reconstruction quality.
-In turn, the generation can also benefit from these improvements, as it is conditioned on $\mathcal{G}_{t}$ whose enhancements could provide more accurate information for generation.
-This reciprocal enhancement motivates an iterative refinement.
-As illustrated in Algorithm 1, we first warm up the optimization for $T_{0}$ steps using conventional reconstruction method [35] on recorded trajectory, and then store the refined video $\mathcal{V}_{t,\text{refine}}^{\prime}$ as a buffer, which will be updated every $K$ steps.
-Due to the additional time demands of running video diffusion model, it is inefficient to set $K=1$, i.e., the buffer is updated every step.
-Please refer to Sec.
-4.3 for the trade-off between different $K$.
-Since the generated contents are conditioned on Gaussian model $\mathcal{G}_{t}$, they are constrained with the recorded scene, thus ensuring the consistency between reconstruction and generation.
-Novel trajectory sampling
-Another important design space is how to sample the novel trajectory $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$.
-The ideal trajectory should balance the following two requirements: first, maximizing the utility of the generative model, i.e., the generated novel views should contribute as much as possible to enhance reconstruction quality; second, minimizing inconsistency between the recorded and generated images.
-Therefore, we adopt a panning camera trajectory that starts from a recorded front view $P^{\prime}_{0}=[R_{0}|T_{0}]$ and gradually shifts laterally:
-$P^{\prime}_{i}=[R_{0}|\frac{i}{F}s\mathbf{v}+T_{0}],$ (10)
-where $\mathbf{v}\in\mathbb{R}^{3}$ is the shifting direction, and $s\in\mathbb{R}$ controls the maximum shifting length.
-This design not only provides novel views beyond the recorded trajectory, but also allows video generation model to extract detailed references from the initial frame.
-At the early stage of optimization, rendering quality may deteriorate as the camera diverges from the recorded trajectory, which offers limited guidance for video generation.
-To address this issue, we initially use a small shifting length $s$ and progressively extend it during optimization, ensuring our method’s stability and robustness across diverse scenes.
-±$\pm 1m$$\pm 2m$$\pm 3m$ Inference speed
+Since the warped image provides a rough overview of the novel trajectory, making it unnecessary to start denoising from scratch when we refine the videos using Eq.
+7.
+Hence, we perturb warped videos by adding Gaussian noise at some noise level.
+This not only reduces unnecessary time costs but also avoids color shifts by preserving accurate low-frequency components in original video.
+Because autonomous vehicles are typically equipped with surround-view cameras, we include side view on recorded trajectory in the same frame as the auxiliary source images, which is also warped using equation (9) to the regions that are invisible by the front view to ensure well-defined unreliability masks for all areas of the novel view.
+Since the sparse depth ground truth in the novel view can be derived from LiDAR measurements, the Gaussian model performs better in depth rendering than in color rendering under novel view, even in the presence of overfitting.
+This allows the warped images to serve as a reliable anchor for indicating the reliability.
+Projected LiDAR as auxiliary condition Additionally, we also utilize colored LiDAR projections as an oracle to guide the generative process.
+For each novel view requiring restoration, LiDAR points projected onto it are accumulated from its adjacent ±$\pm 2$ frames, where points belonging to moving objects are aligned to the rendered frame based on their tracked poses.
+This empirically stabilizes the optimization, especially for those challenging scenes where rendered novel views struggle to provide reliable information in the early optimization.
+In light of the importance of novel view depth supervision within this framework, we additionally use the accumulated LiDAR point cloud as the depth ground truth for novel views.
+Figure 3: Qualitative comparisons on novel trajectories. For each method, images are rendered at the novel views with lateral shift of $3m$ to the recorded trajectories.
+3.4 Training of generative model
+One remaining question is how to obtain a generative prior capable of handling the restoration task in Eq.
+7.
+Due to the limited trajectory in existing driving datasets, precisely simulating degradation in novel trajectory views caused by overfitting has an inherent challenge.
+Fortunately, our inpainting approach can mitigate this issue by disentangling the degraded pattern from the condition, avoiding the necessity to simulate degraded rendering during training and freeing the model from identifying potential degradations.
+Specifically, we employ an edge-aware strategy to generate unreliability masks for the training data.
+This leads to a more general task which can cover scenarios typically encountered during off-trajectory rendering.
+Given a RGB image, we first convert it into grayscale.
+Then its directional derivatives computed using Sobel filters, as the edge intensity at each pixel.
+The resulting intensity map is normalized to yield a probability distribution over the pixels after adding a baseline offset for background regions.
+Then a predefined number of pixels is randomly sampled based on the distribution.
+For each sampled pixel, a $3\times 3$ patch centered at the pixel is masked out to imitate the structured occlusion.
+Both front and side-view camera videos are used in training.
+Because the trajectory of the side-view camera along the moving forward trajectory is similar to the lateral shift of the front cameras adopted in our novel trajectory sampling, as suggested by [28, 4].
+While there remains a gap between training and inference, it does not substantially hinder the usage.
+This is because the training process primarily aims at unleashing the generative capability of the diffusion model to adapt on such a masked image inpainting task, rather than fitting specific mask patterns.
+Actually, even [39] trained only on rendered point clouds can also serve as a competent generative prior for the crafted inpainting-based restoration task.
+While fine-tuning the model on driving videos can further adapt it to the style of driving scenarios, thereby significantly improving its robustness as demonstrated in Sec.
+4.4.
+±$\pm 0m$ (recorded) ±$\pm 1m$$\pm 2m$$\pm 3m$ FPS
 […]
-EmerNeRF [36] 0.1676 0.5149 91.40 0.1372 0.4671 111.67 0.1369 0.4198 132.48 0.12
-$S^{3}$aussian [12] 0.1423 0.5301 70.71 0.0865 0.4781 100.29 0.0430 0.4341 126.22 8.4
-PVG [4] 0.1335 0.5406 71.63 0.0632 0.4686 102.92 0.0365 0.4267 137.12 48
-StreetGaussian [35] 0.2002 0.5945 49.44 0.1668 0.5793 80.05 0.1385 0.5342 105.37 34
-DriveX (Ours) 0.2065 0.5949 48.65 0.1892 0.5827 76.44 0.1755 0.5427 95.60 34
-Table 1: Novel trajectory evaluation on Waymo.
-We compare images rendered in novel trajectories with different shift lengths, where the IoU is computed between the ground truth and those detected by TwinLiteNet [3].
-Once trained, our model solely relies on the Gaussians for inference.
+EmerNeRF [38] 29.60 0.8459 0.1147 0.5101 89.43 0.0881 0.4644 102.98 0.0735 0.4072 122.83 0.12
+$S^{3}$aussian [11] 29.68 0.8670 0.0718 0.4499 115.67 0.0401 0.4127 131.26 0.0177 0.3812 162.03 8.4
+PVG [3] 29.98 0.8627 0.1082 0.5444 51.86 0.0291 0.5102 86.93 0.0170 0.4699 123.53 48
+StreetGaussian [37] 29.76 0.8738 0.2671 0.5919 51.87 0.2122 0.5848 70.66 0.1720 0.5697 93.04 34
+DriveX (ours) 29.74 0.8700 0.2880 0.5932 46.91 0.2710 0.5901 65.52 0.2699 0.5712 79.78 34
+Table 1: Quantitative evaluation on Waymo dataset [25].
+We compare images rendered in novel trajectories with different shift lengths, where the IoU and AP are computed between the projected ground truth and those detected on synthesized views by perception models [21, 2].
 The inference speeds of all methods are measured on the NVIDIA RTX A6000 GPU.
-Gaussian model Lane change Acceleration Deceleration Average
-[…]
-DriveDreamer4D [40] PVG [4] 0.428 53.00 0.411 53.10 0.421 53.78 0.420 53.29
-DriveX (Ours) PVG [4] 0.474 58.53 0.563 65.36 0.611 66.97 0.549 63.62
-DriveX (Ours) StreetGauss. [35] 0.516 62.77 0.569 65.14 0.599 66.23 0.561 64.71
-Table 2: Comparision with DriveDreamer4D [40] under its evaluation setting.
-For fair comparison, we report the same metric proposed in [40] measured on the same sequences, and same Gaussian model [4] with the best reported results in [40].
-3.3 Application on generated videos
-In addition to real-world scenes, the driving videos used for reconstruction can also be created by a video generator [7], which explores new possibilities for the automatic creation of diverse virtual driving environments.
-However, the AI-generated videos present more severe challenges because of the lack of corresponding LiDAR depth and accurate pixel-to-pixel matching between different frames.
-These challenges further necessitate the novel view supervision provided by the generative priors.
-Therefore, we devise an additional recipe for the robust optimization of the AI-generated videos.
-Specifically, since the generated videos cannot be precisely aligned with the conditioned camera parameters, we employ a feed-forward approach [39] to estimate the camera parameters and corresponding depths for each frame.
-For scenes containing moving vehicles, we manually annotate their tracklets as the initialization for the local coordinate frames.
-During optimization, we fine-tune the camera pose to mitigate potential prediction errors.
-In the computation of unreliability, as the generated video is monocular, we use the nearest previous frames that can include the current shifted view content as the auxiliary source images in Eq.
-8 to ensure well-defined unreliability scores for all areas of the novel view.
+Figure 4: Qualitative comparisons with methods incorperating generative model.
+The results of DriveDreamer4D [44] and ReconDreamer [20] are downloaded from their project pages.
+The result of FreeVS [28] is obtained by running its official released code.
+Compared to other alternatives, our method can more effectively harness the generative prior, significantly reducing hallucinations.
 4 Experiments
 4.1 Experimental setup
-Datasets
-For evaluation of the improvement for real-world reconstruction, we conduct experiments on 12 selected sequences comprising surrounding videos and synchronized LiDAR point clouds from Waymo Open Dataset [26].
-Official tracklets are used to crop the point cloud of dynamic traffic participants and initialize their per-frame pose.
-For fair comparison, all images are downsampled into $640\times 920$ in both training and testing.
-To test the applicability of DriveX in generated driving videos, we assess our method on 12 driving scene videos produced by video diffusion models [7].
-Each video is generated given the initial frame from nuScenes validation set as the condition.
-Metrics
-To reasonably benchmark the improvement on the existing reconstruction-based street view synthesis pipeline, we quantitatively compared DriveX with several representative baselines on views from the novel trajectory.
-Since the ground truth sensor data on these trajectories is unavailable, it is infeasible to directly assess the pixel-level fidelity.
+Datasets   We evaluate our framework and train the diffusion model on driving sequences from large-scale Waymo Open Dataset (WOD) [25].
+Each sequence comprises surrounding videos and synchronized LiDAR point clouds capture at 10Hz.
+For training, we use all sequences from its training and validation set.
+The quantitative evaluation is conducted on 18 selected sequences, each containing 40 frames.
+Official bounding boxes are used to crop the point cloud of dynamic traffic participants and initialize their per-frame pose.
+For fair comparison, all images are downsampled into $640\times 960$ in both training and evaluation.
+More details can be found in Appendix B.
+Metrics   We evaluate the view synthesis quality on both recorded and novel trajectories.
+For recorded trajectory, we follow the common practice of reporting the PSNR and SSIM on these views.
+The test views are held out from every 5 frames.
+For novel trajectory views, since the ground truth is unavailable, it is infeasible to directly assess the pixel-level fidelity.
 Therefore, we devised a novel benchmark to comprehensively evaluate the synthesis results at these views.
 Specifically, the Fréchet Inception Distance (FID) [8] between synthesized views in novel trajectories and the captured images from the original trajectories is employed to assess the realism of synthesized images at the distribution level.
-Moreover, we report Lane IoU and vehicle AP to evaluate the distinguishability and fidelity of two safety-critical high-level traffic components, i.e., road lanes and vehicles, in novel trajectory viewpoints.
-Refer to the appendix for detailed calculations of these metrics.
-Implementation details
-For reconstruction of real video where calibrated data and LiDAR are available, we first warm up the optimization with Gaussian model [35] for 50,000 steps and then integrate generative priors for following 30,000 steps.
-For novel trajectory videos, the initial view $P_{0}^{\prime}$ of a novel trajectory $\mathcal{T}^{\prime}$ is selected from every 3 recorded forward cameras, with shifting direction $\mathbf{v}$ oriented to the left or right, and shifting length $s$ ranging from $2m$ to $4m$.
-During the iterative refinement, the buffer for storing refined videos is updated every $K=2000$ steps.
-We use ViewCrafter [37] as our generative prior.
+Moreover, we report Lane IoU and vehicle AP (Refer to Appendix B for these calculations.) to evaluate the distinguishability and fidelity of two safety-critical high-level traffic components, i.e., road lanes and vehicles, in novel trajectory viewpoints.
+Implementation details   For the reconstruction with calibrated camera images and LiDAR point clouds, we first initialize the scene representation [37] by optimizing on the recorded data for $T_{0}=50,000$ steps and then integrate generative priors for following $30,000$ steps.
+For the generation of novel trajectory videos, the initial view $P_{0}^{\prime}$ of a novel trajectory $\mathcal{T}^{\prime}$ is selected from every 3 frames, with shifting length $s$ ranging from $2m$ to $6m$.
+The generative model is initialized from  [39], and then fine-tuned on WOD [25] for $15,000$ iterations with batch size of 24 using 8 RTX A6000 GPUs.
+During the iterative refinement, the buffer for storing refined videos is updated every $K=3000$ steps.
 The refine strength (i.e.
@@
 10 is set to 0.65.
-To alleviate the conflicts between recorded data and generated results on moving cars, their supervision strength is reduced to 0.2.
-For generated videos, we first estimate the camera pose [39] and predict the depth and then perform similar process.
-4.2 Comparison with state of the art
-Results on real-world videos
-For comparisons on real-world reconstruction, we primarily focused on improvements in substantial novel views beyond the recorded trajectory.
-To this end, we synthesized novel driving videos along trajectories shifting from the recorded trajectories by ±$\pm 1m$, ±$\pm 2m$ and ±$\pm 3m$.
-Table 1 demonstrate our significant improvement on FID, AP and IoU, surpassing the leading StreetGaussian [35] with a 9.3% decrease in FID, 1.6% increase in AP and 21.1% increase in IoU under the largest $3m$ shifting length.
+Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement.
+(a) shows the initial restored results without (middle left) and with (middle right) unreliability mask (top right).
+The final rendered results exhibit significant artifacts without mask (bottom left).
+(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right).
+It can be seen that single-time refinement can not completely eliminate artifacts (as shown in orange boxes), resulting in inferior final results compared to iterative refinement (bottom).
+Best viewed with zoom-in.
+4.2 Main results
+In Tab.
+1, we quantitatively compared DriveX with several representative baselines in terms of view synthesis quality and rendering speed.
+For extrapolated novel views, we synthesized driving videos laterally shifting from the recorded trajectories by ±$\pm 1m$, ±$\pm 2m$ and ±$\pm 3m$.
+The result demonstrates our significant improvement on FID, AP and IoU, surpassing the leading reconstruction-based counterpart [37] with a 14.3% decrease in FID and 56.9% increase in IoU under the $3m$ shifting length.
 These enhancements are visually corroborated in Fig.
-3, where safety-critical scene elements can be well synthesized in novel views by our methods.
-It can be observed that by integrating generative priors, our approach demonstrates robust rendering quality under large trajectory deviations, resulting in more realistic driving environments.
-Notably, once trained, our model solely relies on the Gaussians for inference, enabling efficient rendering speed.
-For fair comparison with DriveDreamer4D [40] which is not open-source, we adopt the same experiment setting (sequences, metric and its best reported Gaussian counterpart [4]).
-The results are presented in Tab.
-2 and bottom of Fig.
-3.
-Our model markedly outperforms DriveDreamer4D with variety of Gaussian models on all metrics, visual quality, and the consistency of traffic components.
+3, where our approach demonstrates robust rendering quality under large trajectory deviations, where safety-critical scene elements can be more clearly recognized in novel views by incorporating generative priors.
+Notably, once trained, our model solely relies on the 3DGS for inference, enabling efficient rendering speed.
+For views on recorded trajectories, our method also achieves comparable performance with the baseline.
+The slightly lower metrics often result in perceptually negligible impact on synthesized quality.
+This further indicates our method can seamlessly integrate generative prior by effectively minimizing inconsistencies of generated content.
+Lane change Lane shift ±$\pm 3m$
+[…]
+DriveDreamer4D [44] 0.495 53.42 0.340 51.32
+ReconDreamer [20] 0.554 56.63 0.539 54.58
+DriveX (ours) 0.620 58.60 0.567 58.29
+Lane shift ±$\pm 1m$ Lane shift ±$\pm 3m$
+[…]
+FreeVS [28] 0.2782 0.5777 0.2723 0.5688
+DriveX (ours) 0.2880 0.5932 0.2699 0.5712
+Table 2: Comparison with methods using generative model.
+All results are averaged from 8 sequences used in  [44].
+For lane change, the NTA-IoU and NTL-IoU are evaluated under setting in [44, 20], and we report the best results of [44, 20] from their original paper.
+lane shift is computed in the same way as our main paper.
+For FreeVS [28], we run its official released model and measure the results on the same sequences, note that its cropped image size does not essentially impact the results.
+4.3 Comparison with methods incorporating generative prior
+In previous section, we compared our methods with popular reconstruction-based approaches as the main results because it is directly built upon these reconstruction-based solutions and primarily aims to address the view extrapolation challenge for them.
+Although there are some concurrent works [44, 20, 28, 4, 41] sharing a similar scope with ours, most of them have not released code until our paper submission.
+Despite that, we try our best to compare with those allowing for a reasonable comparison and report the results in Tab.
+2.
+The results indicate that our method performs better than other alternatives, which is also confirmed by qualitative results in Fig.
+4.
+Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints.
+Unlike generative-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than those only relying on sparse LiDAR signal.
+Moreover, our method has significantly faster synthesis speed (34 FPS vs.
+0.43 FPS).
+[…]
+mask all 0.2153 0.6334 102.69
+w/o mask 0.2193 0.6341 76.26
+w/ mask 0.2263 0.6389 74.34
+Table 3: Effects of unreliability mask. “mask all”, “w/o mask”, and “w/ mask” denotes the threshold $\tau$ in Eq. 10 set to 1.0, -1.0, and 0.65, respectively.
+[…]
+500 0.2271 0.6392 75.33 4.8×4.8$\times$4.8 ×
+3,000 0.2263 0.6389 74.34 1.6×1.6$\times$1.6 ×
+6,000 0.2228 0.6388 75.98 1.3×1.3$\times$1.3 ×
+Inf 0.2194 0.6173 89.06 $\mathbf{1.0\times}$
+Table 4: Ablations on buffer intervals.Inf” denotes the novel trajectory supervision only generated one time at iteration $T_{0}$.
+4.4 Ablation study
+Unreliability mask   The unreliability mask is designed to guide the generation of video diffusion by directly indicating where the model needs to be retained or repaired.
+From Tab.
+3 and Fig.
+5(a) we can observe: (i) When all regions are masked, the video diffusion model can only rely on sparse LiDAR projections as condition, leading to inferior results due to the inability to fully leverage information from the in-training Gaussian representation; (ii) When the mask is disabled, the generative model can still refine the rendered results toward the learned video distribution, but usually with too conservative refinement (it also indicates that the model does not strictly follow the given mask during inpainting); (iii) Equipped with the proposed unreliability mask, our full model achieves superior performance in both quantitative and qualitative ablation.
+Iterative refinement   The interval of iterative refinement offers the trade-off between quality and efficiency.
+When the novel trajectory supervision only generated once time at iteration $T_{0}$ (see Fig.
+5(b)), the quality of the novel view is constrained to the in-training Gaussian model, which is overfitted on the recorded trajectory.
+To the other extreme, we are unable to afford the time cost of running the diffusion at every step.
+From Tab.
+4, it can be seen that even a small interval 500 leads to 4.8×4.8$\times$4.8 × training time.
+Hence we finally set the interval to 3,000 to strike the balance.
+Figure 6: Domain-specific fine-tuning enhances the style adaptation on driving videos.
+Here the appearance of the black car refined by model without fine-tuning seems to lack details and has unrealistic material.
+Projected LiDAR points as auxiliary oracle   Although using the masked rendered image as the only condition is sufficient to achieve satisfactory performance in most scenes, incorporating projected LiDAR points as the additional oracle can greatly improve robustness in some difficult cases.
+In these scenes, initial poor novel view renderings would result in large masked regions, LiDAR projection can provide direct guidance for precisely restoring them.
+This resulted enhancement is reflected by the quantitative comparison in Tab.
+5.
+[…]
+render LiDAR ft.
+0.2092 0.6257 90.30
+0.2123 0.6334 86.30
+0.2153 0.6225 75.18
+0.2263 0.6389 74.34
+Table 5: Ablations on the components of generative model.
+“w/ render” means using masked novel view image as condition.
+“w/ LiDAR” denotes adopting LiDAR projection as condition.
+“w/ ft.” represents training diffusion model on driving video data [25].
+Fine-tuning generative model on driving videos   The comparison between the $2^{nd}$ and $4^{th}$ row of Tab.
+5 demonstrates that the diffusion model fine-tuned on in-domain driving videos outperforms that trained solely on general-purpose 3D datasets [39], especially in FID.
+This improvement comes from the generative model’s adaptation to the style of real driving scenes.
+As shown in Fig.
+6, the counterpart without domain-specific fine-tuning tends to generate cartooned foreground vehicles without detailed texture.
+Therefore, beyond the improvement in quantitative metrics, it also simplifies our pipeline by removing the need for some hand-designed processing, such as lowering the refine weight for foreground vehicles.
+5 Conclusion
+In this paper, we present a novel framework that integrates generative priors into the reconstruction of the driving scene from single-trajectory recorded videos, addressing limitations in extrapolating novel views beyond a recorded trajectory.
+To reasonably leverage the video generative model to provide consistent novel view supervision, we construct a tailored inpainting-based video restoration task to refine the novel view renderings for optimizing the street scene representation.
+The experiments demonstrate that the proposed method significantly improves the synthesis quality for novel trajectories.
+These results pave the new way for driving scene synthesis on free-form trajectories.
+Acknowledgments
+This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).
+Appendix
+Appendix A Discussion on limitations
+Beyond demonstrating the efficacy of the proposed method in synthesizing driving views along arbitrary trajectories, this paper also acknowledges several limitations.
+First, incorporating a diffusion-based generative model significantly increases reconstruction time.
+But such complexity might be unnecessary, as the required generative diversity is substantially constrained under our inpainting formulation.
+suggesting that a more lightweight model can be employed in the future.
+Second, our method currently employs a simple panning camera scheme to sample novel trajectories to be restored.
+However, this fix design may not be optimal for all scenarios, and leave some hand-designed components in the pipeline.
+Future work could consider to adaptively explore the optimal novel trajectories in optimization.
+Appendix B More implementation details
+Metric details
+Previous driving scene reconstruction methods [40, 7, 12] typically utilize PSNR, SSIM [31] and LPIPS [43] for evaluate image level similarity with ground truth.
+However, these metrics are inapplicable for evaluating free-form trajectories due to the absence of corresponding ground-truth images for novel trajectories.
+Therefore, inspired by [28, 44, 20], we propose a novel set of metrics to evaluate the recognizability of high-level traffic elements in synthesized views, based on ground truth annotations in 3D space.
+For vehicles, we projected their ground truth 3D bounding boxes onto the novel views to obtain corresponding 2D bounding boxes.
+Since IoU lacks the precision granularity needed to understand performance across various intersection over union threshold.
+Therefore, we calculate the Average Precision (AP[50:95]) between the projected ground truth and those detected by Faster R-CNN [21], which provides a more nuanced assessment than IoU.
+For road lanes, NTL-IoU adopted in  [44, 20] calculates the mIoU between the lanes on synthesized views detected by [2] and the ground truth, averaging both foreground lanes and backgrounds.
+However, including the background can inflate accuracy scores due to the easy detection of non-lane areas.
+Therefore, following standard lane segmentation evaluation [2], we focus specifically on IoU with only the foreground to better evaluate lane synthesis quality.
+Warping operation $\phi$ in Eq. (8) of main paper
+Given a camera extrinsic matrix $E\in\mathbb{R}^{4\times 4}$ and intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, the mapping from world points $p_{\text{world}}\in\mathbb{R}^{3}$ to pixel coordinates $(x,y)$ and depth $d$ can be expressed by:
+$\begin{bmatrix}p_{\text{camera}}\\ 1\end{bmatrix}$ $=(*,*,d,1)^{\top}=E\begin{bmatrix}p_{\text{world}}\\ 1\end{bmatrix}$ (11)
+$(x_{*},y_{*},d)^{\top}$ $=Kp_{\text{camera}}$ (12)
+$(x,y)$ $=(\frac{x_{*}}{d},\frac{y_{*}}{d})$ (13)
+Note that all the calculation above is invertible, allowing the mapping from $(x,y,d)$ to world points $p_{\text{world}}$.
+Validation sequences
+There are 18 curated driving sequences from WOD [25] used for qualitative and quantitative evaluations in our paper. The names of all segments are listed below:
+  • 10359308928573410754_720_000_740_000
+  • 11450298750351730790_1431_750_1451_750
+  • 12496433400137459534_120_000_140_000
+  • 15021599536622641101_556_150_576_150
+  • 16767575238225610271_5185_000_5205_000
+  • 17860546506509760757_6040_000_6060_000
+  • 3015436519694987712_1300_000_1320_000
+  • 6637600600814023975_2235_000_2255_000
+  • 10444454289801298640_4360_000_4380_000
+  • 10588771936253546636_2300_000_2320_000
+  • 10625026498155904401_200_000_220_000
+  • 11017034898130016754_697_830_717_830
+  • 1191788760630624072_3880_000_3900_000
+  • 11928449532664718059_1200_000_1220_000
+  • 14810689888487451189_720_000_740_000
+  • 4414235478445376689_2020_000_2040_000
+  • 6242822583398487496_73_000_93_000
+  • 7670103006580549715_360_000_380_000
+Unless stated otherwise, all quantitative ablation experiments are conducted on segments 15021599536622641101_556_150_576_150 and 3015436519694987712_1300_000_1320_000, while the results are evaluated on novel trajectories with lateral shift $3m$.
+Appendix C Additional ablations
+[…]
+0.4 0.2184 0.6202 96.21 $\mathbf{1.0\times}$
+0.6 0.2263 0.6389 74.34 $\times1$
+0.8 0.2192 0.6259 81.55 1.2×1.2$\times$1.2 ×
+1.0 0.2113 0.6293 106.01 1.3×1.3$\times$1.3 ×
+Table 6: Effect of the noise level. The noise level $l$ denotes adding the noise corresponding to the timestep $t=\lfloor lN\rfloor$, where $N$ is the total step needed for denoising from scratch.
+[…]
+PVG [3] 0.0200 0.5401 118.41
+PVG [3] 0.1872 0.6197 91.85
+StreetGaussian [37] 0.1217 0.6124 103.08
+StreetGaussian [37] 0.2263 0.6389 74.34
+Table 7: Ablation study on different Gaussian model.
+[…]
+w/o progress. 0.2204 0.6372 75.41
+w/o warp 0.2196 0.6357 105.77
+Full model 0.2263 0.6389 74.34
+Table 8: Ablations on other design choices.
+“w/o progress” denotes that the shift length is fixed at $4m$ throughout the optimization.
+“w/o warp” means that the denoising is started from the noisy rendered images rather than warped views.
+Refine strength
+When we refine the rendered videos using video diffusion, we also study the effect the refine strength (level of noise added to the images).
+In Tab.
+6, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time.
+With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement.
+On the other hand, a high refine strength yields better novel views but may not keep fidelity to the original scene.
+Additionally, higher refine strength requires more denoising steps, increasing the training time.
+Consequently, refine strength of 0.6 is found to be the best balance.
+Different Gaussian model
+Although we adopt static 3D Gaussian splatting [37] as the default scene representation in previous experiments, our method is also capable of synergy with different choices.
+To demonstrate the versatility of DriveX, we also evaluate our method on a representative dynamic Gaussian representation for driving view synthesis [3].
+From Tab.
+7, we can observe that our strategy consistently reduces the FID by a large margin by substantially reducing artifacts in novel views, and effectively recovering street lanes and cars (i.e.
+IoU, AP) that are entirely indiscernible in the baseline.
 w/o ours
@@
 The first and the second rows show the novel trajectory synthesis results by the baseline [37] and our method, respectively.
+Other designs
+We also ablate other design choices in Tab.
+8, including progressive increasing shifted length and warped condition.
+The results show that the final quality degrades without any element, which indicates their helpfulness on the rendering quality and the stability of optimization.
 Method FID $\downarrow$
@@
 Table 9: Quantitative comparison on generated video. We compared the FID score on videos generated by Vista [6].
-Results on generated videos
+Appendix D Comparison with methods incorporating generative prior
+Since our work is directly built upon the reconstruction-based driving simulation methods and primarily aims to address the view extrapolation challenge within them, the comparisons in the main paper are mainly against the reconstruction-based approaches.
+While we acknowledge some concurrent works [44, 20, 28, 4, 41] are sharing a similar scope with ours, most of them have not released their code, making comprehensive and rigorous comparison with them infeasible.
+Despite that, we try our best to compare with those methods that allow for a reasonable comparison and report the results in Tab.
+2.
+The results indicate that our method performs better than other alternatives, which is also confirmed by the qualitative results in Fig.
+4.
+Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints.
+Compared to generation-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than that only relying on sparse LiDAR signal.
+Moreover, our method has significantly faster synthesis speed (34 FPS vs.
+0.43 FPS).
+Appendix E Application on generated videos
 Another intriguing application of the proposed method is transforming AI-generated video into a re-enactable driving world.
-To quantitatively evaluate this capability, we deploy DriveX and the reconstruction-only baseline [35] on 12 driving videos generated by [7].
+However, the AI-generated videos present more severe challenges because of the lack of corresponding LiDAR depth and accurate pixel-to-pixel matching between different frames.
+These challenges further necessitate the novel view supervision provided by the generative priors.
+Therefore, we devise an additional recipe for the robust optimization of the AI-generated videos.
+Details
+Specifically, since the generated videos cannot be precisely aligned with the conditioned camera parameters, we employ a feed-forward approach [42] to estimate the camera parameters and corresponding depths for each frame.
+For scenes containing moving vehicles, we manually annotate their tracklets as the initialization for the local coordinate frames.
+During optimization, we fine-tune the camera pose to mitigate potential prediction errors.
+In the computation of unreliability, as the generated video is monocular, we use the nearest previous frames that can include the current shifted view content as the auxiliary source images in Eq.
+(9) of main paper to ensure well-defined unreliability scores for all areas of the novel view.
+Besides, we include an entropy loss for the opacity of each Gaussian to prevent potential overfitting.
+Results
+To quantitatively evaluate this capability, we compared DriveX and the reconstruction-only baseline [37] on 12 driving videos generated by Vista [6].
+Due to the evident stylistic gap between the generated videos and those in WOD, we only use the diffusion model trained on general-purpose 3D datasets as generative prior without LiDAR projection as condition.
 The FID of videos rendered across different novel trajectories is reported in Tab.
 9.
-The results indicate that the realism of rendered videos is well preserved within a certain range of shifted length.
+The results indicate that the realism of rendered videos is effectively preserved within a certain range of shifted length.
 Compared to the reconstruction-only baseline, our DriveX achieves 4.1% and 2.6% FID reduction in shifting length of ±$\pm 1m$ and ±$\pm 2m$ respectively.
-To obtain real world distance, we align the predicted depth and LiDAR depth of the first frame by the least square estimation.
+To obtain real-world distance, we align the predicted depth and LiDAR depth of the first frame by the least square estimation.
 These advancements are also clearly illustrated in Fig.
-4.
+7.
 As can be seen, the baseline method exhibits significant degradation in views that are far from the recorded trajectory.
-4.3 Ablation study
-[…]
-PVG [4] 0.0200 0.5401 118.41
-PVG [4] 0.0819 0.6188 101.04
-StreetGaussian [35] 0.1217 0.6124 103.08
-StreetGaussian [35] 0.2092 0.6257 90.30
-Table 4: Ablation study on different Gaussian model. Unless stated otherwise, all ablation experiments are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
-[…]
-mask all 0.0899 0.5781 107.58
-w/o mask 0.1106 0.5965 90.49
-w/ mask 0.2092 0.6257 90.30
-Table 5: Effects of unreliability mask.
-[…]
-500 0.1070 0.5954 86.56 4.8×4.8$\times$4.8 ×
-2,000 0.2092 0.6257 90.30 1.7×1.7$\times$1.7 ×
-5,000 0.1104 0.5992 105.25 1.4×1.4$\times$1.4 ×
-Inf 0.0945 0.5845 95.08 1.0×1.0$\times$1.0 ×
-Table 6: Ablations on buffer intervals. Inf denotes the novel trajectory supervision only generated once time at iteration $T_{0}$.
-(a) Ablation on unreliability mask (b) Ablation on iterative refinement
-Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement.
-Fig.(a) shows the refined results without (middle left) and with (middle right) unreliability mask (top right).
-The final rendered results exhibit significant artifacts without mask (bottom left).
-Fig.(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right).
-It can be seen that single-time refinement can not completely eliminate artifacts, resulting in inferior final rendered results compared to iterative refinement (bottom).
-$T_{0}$ and $T$ indicate the first and last iteration after warm up, respectively.
-Different Gaussian model
-Although we adopt StreetGaussian [35] as the default Gaussian counterpart in previous experiments, our method is also capable of synergy with different choices.
-To demonstrate the versatility of DriveX, we evaluate on a dynamic counterpart PVG [4].
-From Tab.
-4, we can observe that our strategy consistently reduces the FID by a large margin, and effectively recovering street lanes and cars (i.e.
-IoU, AP) that are entirely indiscernible in the baseline methods.
-Unless stated otherwise, all ablation experiments are conducted on two Waymo sequences.
-Unreliability mask
-The unreliability mask is designed to guide the generation of video diffusion.
-From Tab.
-5 and Fig.
-5(a) we can observe: (i) When all regions are masked, where video diffusion model does not know any information from the in-training Gaussian, leading to inferior results; (ii) When the mask is disabled, the generative model can still refine the rendered results toward the video distribution, but with minor refinement; (iii) Equipped with proposed unreliability mask, our full model achieves superior performance in both quantitative and qualitative ablation.
-Iterative refinement
-The interval of iterative refinement offers the trade-off between quality and efficiency.
-When the novel trajectory supervision only generated at iteration $T_{0}$ (see Fig.
-5(b)), the quality of the novel view is constrained to the in-training Gaussian model which is not fully optimized.
-To the other extreme, we are unable to afford the time cost of running the diffusion at every step.
-From Tab.
-6, a small interval 500 leads to 4.8×4.8$\times$4.8 × training time.
-Hence we finally chose the interval 2,000 to strike the balance.
-5 Conclusion
-In this paper, we present a novel framework that integrates generative priors into the reconstruction of the driving scene from single-trajectory recorded videos, addressing limitations in extrapolating novel views beyond a recorded trajectory.
-To reasonably leverage the video generative model to provide consistent novel view supervision, we construct a tailored inverse problem to refine the novel view rendered images and progressively enhance the street scene model.
-The extensive experiments demonstrate that the proposed method significantly improves the synthesis quality for novel trajectories, both in real and AI-generated scenes.
-These results pave the new way for driving scene synthesis on free-form trajectories.
-Supplementary Material
-6 More implementation details
-Used sequences
-The results in Table 1 of the main text are computed from 12 sequences curated from WOD [26]. The names of all sequences are listed below:
-  • segment-10444454289801298640
-  • segment-10625026498155904401
-  • segment-6242822583398487496
-  • segment-8822503619482926605
-  • segment-10359308928573410754
-  • segment-11450298750351730790
-  • segment-12496433400137459534
-  • segment-15021599536622641101
-  • segment-16767575238225610271
-  • segment-17860546506509760757
-  • segment-3015436519694987712
-  • segment-6637600600814023975
-For the reconstruction from generated videos, the videos are generated given the following images from nuScenes [2] validation set:
-  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537853021662460
-  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984433512460
-  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537852899162460
-  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984912912460
-  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984834412460
-  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984833912460
-  • n015-2018-07-18-11-41-49+0800__CAM_FRONT__1531885347362460
-  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984560512460
-  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984563012460
-  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657697012404
-  • n008-2018-09-18-14-35-12-0400__CAM_FRONT__1537295844362404
-  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657694012404
-Metric details
-Previous driving scene reconstruction methods [38, 8, 13] typically utilize PSNR, LIPIS and SSIM [31] for quality evaluation.
-However, these metrics are inapplicable for evaluating free-form trajectories due to the absence of corresponding ground-truth images for novel trajectories.
-Latest research [40] proposed Novel Trajectory Agent IoU (NTA-IoU) and Novel Trajectory Lane IoU (NTL-IoU) for traffic components evaluation.
-Accordingly, we compared NTA-IoU and NTL-IoU, as shown in Table 2.
-Furthermore, we propose improvements to the evaluation metrics of NTA-IoU and NTL-IoU.
-For vehicle quality, similar to NTA-IoU, we projected the original 3D vehicle bounding boxes onto the new viewpoints to generate corresponding 2D bounding boxes.
-IoU lacks the precision granularity needed to understand performance across various intersection over union threshold.
-Therefore, we calculate the Average Precision (AP[50:95]) which provides a more nuanced assessment of detection accuracy.
-For lane quality, NTL-IoU calculates the mIoU between the rendered and the ground truth lanes, averaging both foreground lanes and backgrounds.
-However, including the background can inflate accuracy scores due to the easy detection of non-lane areas.
-Therefore, following official lane detection benchmarks [3] we focus specifically on IoU with only the foreground to better evaluate lane recognition performance.
-For the evaluation of the capability as driving the scene generator, we mainly assess the realism of synthesized images across different trajectories by measuring their FID with the images at recorded trajectories.
-More details
-For reconstruction from real data, we adopt regularization loss as introduced in [35], including sky loss, LiDAR depth loss and entropy loss for objects.
-In our framework, the LiDAR depth loss is used for both recorded trajectory and novel trajectory.
-In addition, we balance the recorded and generated data with 50% of the training data being generated.
-For reconstruction from generated videos, we additionally include an entropy loss for the opacity of each Gaussian to prevent potential overfitting.
-Warping operation $\phi$ in Eq. 7
-Given a camera extrinsic matrix $E\in\mathbb{R}^{4\times 4}$ and intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, the mapping from world points $p_{\text{world}}\in\mathbb{R}^{3}$ to pixel coordinates $(x,y)$ and depth $d$ can be expressed by:
-$\begin{bmatrix}p_{\text{camera}}\\ 1\end{bmatrix}$ $=(*,*,d,1)^{\top}=E\begin{bmatrix}p_{\text{world}}\\ 1\end{bmatrix}$ (11)
-$(x_{*},y_{*},d)^{\top}$ $=Kp_{\text{camera}}$ (12)
-$(x,y)$ $=(\frac{x_{*}}{d},\frac{y_{*}}{d})$ (13)
-Note that all the calculation above is invertible, allowing the mapping from $(x,y,d)$ to world points $p_{\text{world}}$.
-7 More ablations
-Refine strength
-[…]
-0.4 0.1122 0.5904 98.31 1.0×1.0$\times$1.0 ×
-0.6 0.2092 0.6257 90.30 $\times1$
-0.8 0.0996 0.5874 92.26 1.2×1.2$\times$1.2 ×
-1.0 0.0773 0.5979 117.91 1.3×1.3$\times$1.3 ×
-Table 7: Effect of the noise level.
-In this table, we evaluate the performance under different refine strengths.
-The metrics are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
-When we refine the rendered videos using video diffusion, we also study the effect the refine strength (level of noise added to the images).
-In Tab.
-7, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time.
-With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement.
-On the other hand, a high refine strength yields better novel views but may not keep fidelity to the original scene.
-Additionally, higher refine strength requires more denoising steps, increasing the training time.
-Consequently, a refine strength of 0.6 is found to be the best balance.
-Other designs
-[…]
-w/o pose refinement N/A N/A 116.86
-w/o depth prior N/A N/A 98.62
-Full model N/A N/A 95.73
-w/o progress. 0.1075 0.5959 91.22
-w/o warp 0.0652 0.5962 119.24
-Full model 0.2092 0.6257 90.30
-Table 8: Ablations on other design choices. The metric in the first three lines is measured on the generated video, the others are tested on real scene.
-We also ablate other design choices in Tab.
-8, including the pose refinement and depth prior in the reconstruction from generative video, as well as progressive increasing shifted length and warped condition.
-The results show that the final quality degrades without any element.
-The results indicate that the exclusion of any element leads to a decline in the rendering quality and the stability of optimization.
-8 Comparison with video generative model
-FID $\downarrow$
-±$\pm 0m$$\pm 1m$$\pm 2m$
-MagicDrive-t [6] 125.07 130.23 138.31
-Vista [7] 97.48 99.19 148.07
-Recon-only [35] 41.55 95.32 144.01
-Ours 36.54 71.47 119.11
-Table 9: Comparison with the video-based driving world model.
-The FID score is calculated between the recorded videos from the nuScenes dataset [2] and the novel trajectory videos synthesized by each method.
-In this section, we compare the novel trajectory video synthesis with representative open-sourced generative-based driving world models [7, 6].
-Since these works are primarily trained on nuScenes dataset [2], we still conduct the experiments on 12 sequences from this dataset mentioned in Sec.
-6.
-Different from the experiment in Tab.
-3, in this experiment, our method takes ground truth recorded trajectory videos as input.
-The FID between the recorded videos and the synthesized novel trajectory videos from the different methods are reported in Tab.
-9.
-Note that it is difficult to precisely control the trajectory for generative model.
-Though we adjust the action control, it still tend to generate videos in the original trajectory.
-Although all methods do not have significant artifacts, DriveX still demonstrates a clear advantage in FID metrics.
-This is because it can effectively avoid hallucinations that may be produced by generative models, making its result better aligned with the ground truth.
-Another advantage difficult to demonstrate by this metric is that DriveX can achieve precise camera pose control.
-In contrast, other methods are constrained by the diversity and imbalance of actions existing in public driving datasets, thus making it difficult to achieve accurate free-form trajectory synthesis.
-9 More results
-Figure 6: More comparison results.
-Fig. 6 shows more comparison results. We recommend readers to watch the video in the supplementary material.
+Figure 8: Results on the PandaSet dataset [35]. The results are on the novel trajectory shifted $2.0m$ from the recoded trajectory.
+Appendix F Results on the PandaSet dataset
+To further validate the versatility of the proposed method on a broader range of scene types and sensor configurations, we apply our approach to two sequences indexed $001$ and $003$ from the PandaSet dataset [35], each consisting of 80 frames.
+The image resolution is downsampled to $450\times 800$ for both training and evaluation.
+In Fig.
+8, we compared our method with the state-of-the-art reconstruction baseline StreetGaussian [37] on the trajectory with shift length $2.0m$.
+The result suggests that our method can significantly reduce artifacts in synthesized images and improve the discernibility of safety-critical traffic elements, which is also corroborated by our better FID.
+Appendix G More quanlitative results
+Figure 9: Results on more aggressive view change.
+Figure 10: Failure cases. cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.
+More aggressive view change.
+We test the proposed method on some challenging cases to assess its robustness.
+In Fig.
+9 (left), the view is rendered from the camera elevated by $3m$ above the recorded trajectory, with a pitch angle of $30^{\circ}$.
+In Fig.
+9 (right), we increase the lateral shift length to $6m$.
+To accommodate these aggressively changed views, we include them in the refined trajectory during optimization.
+As shown in the results, the baseline method almost totally fails to synthesize meaningful images under these viewpoints, while our solution substantially improves the fidelity.
+Failure case.
+Despite the robustness and generalization ability on diverse scenes as demonstrated above, it should be noted that the proposed framework only aims to provide regularization for unconstrained reconstruction.
+As a reconstruction pipeline, it does not work well on totally unseen regions of the scene.
+For example, in Fig.
+10, our method cannot create a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views and thus lacks proper geometry initialization before the refinement stage.
+Figure 11: The evolution of unreliability mask during optimization.
+The evolution of unreliability mask during optimization.
+In Fig.
+11, we present novel view renderings and the corresponding unreliability masks at different training iterations.
+The results clearly show that the reliable regions gradually expand during iterative refinement, demonstrating the effectiveness of the proposed unreliability mask mechanism–it can serve as a reliable cue for the rendered images rather than introducing compounding errors.
+Additionally, the distribution of the mask indicates that unreliable regions mainly appear near edge areas of the visual elements, which is the rationale behind the adoption of edge-aware inpainting.
+More comparisons with sota methods The video comparison with the state-of-the-art alternates can be found in our project page.

### v3

@@
-DriveX: Driving View Synthesis on Free-form Trajectories with Generative Prior
+Driving View Synthesis on Free-form Trajectories with Generative Prior
 Abstract
@@
 To further enhance the consistency and fidelity of generated contents, the pseudo ground truth is progressively updated with gradually improved novel trajectory rendering, allowing both components to co-adapt and reinforce each other while minimizing the disruption on the optimization.
-By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time—unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
+By tightly integrating 3D scene representation with generative prior, DriveX achieves high-quality view synthesis beyond recorded trajectories in real time–unlocking new possibilities for flexible and realistic driving simulations on free-form trajectories.
 Figure 1: Though conventional reconstruction-based approaches can well fit views on the recorded trajectory, they struggle to extrapolate onto the views beyond it (top).
@@
 10 is set to 0.65.
-Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement.
-(a) shows the initial restored results without (middle left) and with (middle right) unreliability mask (top right).
-The final rendered results exhibit significant artifacts without mask (bottom left).
-(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right).
-It can be seen that single-time refinement can not completely eliminate artifacts (as shown in orange boxes), resulting in inferior final results compared to iterative refinement (bottom).
-Best viewed with zoom-in.
 4.2 Main results
@@
 This further indicates our method can seamlessly integrate generative prior by effectively minimizing inconsistencies of generated content.
+Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement.
+(a) shows the initial restored results without (middle left) and with (middle right) unreliability mask (top right).
+The final rendered results exhibit significant artifacts without mask (bottom left).
+(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right).
+It can be seen that single-time refinement can not completely eliminate artifacts (as shown in orange boxes), resulting in inferior final results compared to iterative refinement (bottom).
+Best viewed with zoom-in.
 Lane change Lane shift ±$\pm 3m$
@@
 Hence we finally set the interval to 3,000 to strike the balance.
-Figure 6: Domain-specific fine-tuning enhances the style adaptation on driving videos.
-Here the appearance of the black car refined by model without fine-tuning seems to lack details and has unrealistic material.
-Projected LiDAR points as auxiliary oracle   Although using the masked rendered image as the only condition is sufficient to achieve satisfactory performance in most scenes, incorporating projected LiDAR points as the additional oracle can greatly improve robustness in some difficult cases.
-In these scenes, initial poor novel view renderings would result in large masked regions, LiDAR projection can provide direct guidance for precisely restoring them.
-This resulted enhancement is reflected by the quantitative comparison in Tab.
-5.
-[…]
-render LiDAR ft.
-0.2092 0.6257 90.30
-0.2123 0.6334 86.30
-0.2153 0.6225 75.18
-0.2263 0.6389 74.34
-Table 5: Ablations on the components of generative model.
-“w/ render” means using masked novel view image as condition.
-“w/ LiDAR” denotes adopting LiDAR projection as condition.
-“w/ ft.” represents training diffusion model on driving video data [25].
-Fine-tuning generative model on driving videos   The comparison between the $2^{nd}$ and $4^{th}$ row of Tab.
-5 demonstrates that the diffusion model fine-tuned on in-domain driving videos outperforms that trained solely on general-purpose 3D datasets [39], especially in FID.
-This improvement comes from the generative model’s adaptation to the style of real driving scenes.
-As shown in Fig.
-6, the counterpart without domain-specific fine-tuning tends to generate cartooned foreground vehicles without detailed texture.
-Therefore, beyond the improvement in quantitative metrics, it also simplifies our pipeline by removing the need for some hand-designed processing, such as lowering the refine weight for foreground vehicles.
 5 Conclusion
@@
 In Tab.
-6, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time.
+5, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time.
 With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement.
@@
 From Tab.
-7, we can observe that our strategy consistently reduces the FID by a large margin by substantially reducing artifacts in novel views, and effectively recovering street lanes and cars (i.e.
+6, we can observe that our strategy consistently reduces the FID by a large margin by substantially reducing artifacts in novel views, and effectively recovering street lanes and cars (i.e.
 IoU, AP) that are entirely indiscernible in the baseline.
+Figure 6: Domain-specific fine-tuning enhances the style adaptation on driving videos.
+Here the appearance of the black car refined by model without fine-tuning seems to lack details and has unrealistic material.
+Projected LiDAR points as auxiliary oracle   Although using the masked rendered image as the only condition is sufficient to achieve satisfactory performance in most scenes, incorporating projected LiDAR points as the additional oracle can greatly improve robustness in some difficult cases.
+In these scenes, initial poor novel view renderings would result in large masked regions, LiDAR projection can provide direct guidance for precisely restoring them.
+This resulted enhancement is reflected by the quantitative comparison in Tab.
+8.
 w/o ours
@@
 The first and the second rows show the novel trajectory synthesis results by the baseline [37] and our method, respectively.
+[…]
+render LiDAR ft.
+0.2092 0.6257 90.30
+0.2123 0.6334 86.30
+0.2153 0.6225 75.18
+0.2263 0.6389 74.34
+Table 8: Ablations on the components of generative model.
+“w/ render” means using masked novel view image as condition.
+“w/ LiDAR” denotes adopting LiDAR projection as condition.
+“w/ ft.” represents training diffusion model on driving video data [25].
+Fine-tuning generative model on driving videos   The comparison between the $2^{nd}$ and $4^{th}$ row of Tab.
+8 demonstrates that the diffusion model fine-tuned on in-domain driving videos outperforms that trained solely on general-purpose 3D datasets [39], especially in FID.
+This improvement comes from the generative model’s adaptation to the style of real driving scenes.
+As shown in Fig.
+6, the counterpart without domain-specific fine-tuning tends to generate cartooned foreground vehicles without detailed texture.
+Therefore, beyond the improvement in quantitative metrics, it also simplifies our pipeline by removing the need for some hand-designed processing, such as lowering the refine weight for foreground vehicles.
 Other designs
 We also ablate other design choices in Tab.
-8, including progressive increasing shifted length and warped condition.
+7, including progressive increasing shifted length and warped condition.
 The results show that the final quality degrades without any element, which indicates their helpfulness on the rendering quality and the stability of optimization.
@@
 Table 9: Quantitative comparison on generated video. We compared the FID score on videos generated by Vista [6].
-Appendix D Comparison with methods incorporating generative prior
-Since our work is directly built upon the reconstruction-based driving simulation methods and primarily aims to address the view extrapolation challenge within them, the comparisons in the main paper are mainly against the reconstruction-based approaches.
-While we acknowledge some concurrent works [44, 20, 28, 4, 41] are sharing a similar scope with ours, most of them have not released their code, making comprehensive and rigorous comparison with them infeasible.
-Despite that, we try our best to compare with those methods that allow for a reasonable comparison and report the results in Tab.
-2.
-The results indicate that our method performs better than other alternatives, which is also confirmed by the qualitative results in Fig.
-4.
-Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints.
-Compared to generation-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than that only relying on sparse LiDAR signal.
-Moreover, our method has significantly faster synthesis speed (34 FPS vs.
-0.43 FPS).
-Appendix E Application on generated videos
+Appendix D Application on generated videos
 Another intriguing application of the proposed method is transforming AI-generated video into a re-enactable driving world.
@@
 Figure 8: Results on the PandaSet dataset [35]. The results are on the novel trajectory shifted $2.0m$ from the recoded trajectory.
-Appendix F Results on the PandaSet dataset
+Appendix E Results on the PandaSet dataset
 To further validate the versatility of the proposed method on a broader range of scene types and sensor configurations, we apply our approach to two sequences indexed $001$ and $003$ from the PandaSet dataset [35], each consisting of 80 frames.
@@
 The result suggests that our method can significantly reduce artifacts in synthesized images and improve the discernibility of safety-critical traffic elements, which is also corroborated by our better FID.
-Appendix G More quanlitative results
+Appendix F More quanlitative results
 Figure 9: Results on more aggressive view change.
-Figure 10: Failure cases. cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.
+Figure 10: Failure cases. Our method cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.
 More aggressive view change.